### v1

# SynthesizRR: Generating Diverse Datasets with Retrieval Augmentation

Abstract
Large language models (LLMs) are versatile and can address many tasks, but for computational efficiency, it is often desirable to distill their capabilities into smaller student models. One way to do this for classification tasks is via dataset synthesis, which can be accomplished by generating examples of each label from the LLM. Prior approaches to synthesis use few-shot prompting, which relies on the LLM’s parametric knowledge to generate usable examples. However, this leads to issues of repetition, bias towards popular entities, and stylistic differences from human text. In this work, we propose Synthesize by Retrieval and Refinement (SynthesizRR), which uses retrieval augmentation to introduce variety into the dataset synthesis process: as retrieved passages vary, the LLM is “seeded” with different content to generate its examples. We empirically study the synthesis of six datasets, covering topic classification, sentiment analysis, tone detection, and humor, requiring complex synthesis strategies. We find SynthesizRR greatly improves lexical and semantic diversity, similarity to human-written text, and distillation performance, when compared to standard 32-shot prompting and six baseline approaches.
SynthesizRR: Generating Diverse Datasets with Retrieval Augmentation
[…]
1 Introduction
Figure 1: Synthetic examples from few-shot generation (middle) and SynthesizRR (bottom). Few-shot generation leverages a frozen LLM’s parametric knowledge as the source of content in the generated examples. SynthesizRR incorporates a distinct content sourcing step by retrieving documents from a corpus. For the task of detecting political bias, a news article is retrieved and transformed in tone. The resulting synthesis procedure yields diverse examples which more closely match human-written examples (top).
Figure 2: Abstract depiction of the SynthesizRR pipeline. In the content sourcing stage, we retrieve $K$ unique document $\{r_{1},\dots,r_{K}\}$ from a large corpus for each in-context covariate $x_{\text{ICL}}$. The task-inversion stage of synthesis uses a parameterized context refinement prompt $\mathcal{P}_{\tau}$, which takes parameters $R_{inv}$ (inversion instruction), $r_{k}$ (a retrieved document), and $\mathcal{V}(y_{\text{ICL}})$ (the verbalized target label). A generalist teacher LLM autoregressively generates a synthetic covariate. Each in-context example thus produces $K$ unique synthetic examples $\{\tilde{x}_{1},\dots,\tilde{x}_{K}\}$, which we include in the dataset with target $y_{\text{ICL}}$.
Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023; Bubeck et al., 2023), LLaMa (Touvron et al., 2023) and Claude (Bai et al., 2022) are powerful generalist models, capable of solving multiple tasks without parameter tuning via zero-shot or few-shot prompting. In comparison, previous approaches fine-tuned variants of BERT (Devlin et al., 2019) on task-specific demonstrations, producing specialist models. These smaller specialist models are more economical at inference time, but typically require at least thousands of examples to train.
Recent work has sought to avoid this reliance on annotation by fine-tuning specialist models on synthetic datasets via teacher-student distillation (West et al., 2022). While this has been employed in the context of classification (Yu et al., 2023a; Ye et al., 2022a, b), human-preference alignment (Lee et al., 2023; Bai et al., 2022), language understanding (Meng et al., 2022; Schick and Schütze, 2021), and even tabular data (Borisov et al., 2022), synthetic data has limitations. Yu et al. (2023a) examine the data-generation process and discover that naive prompts generate text inputs that are low in diversity and regurgitate the biases of the teacher LLMs, which proliferate into the student model.
Figure 1 illustrates this: few-shot generation (Brown et al., 2020) is a common approach that uses an LLM to perform a multitude of tasks without modifying its parameters. Prior work has used few-shot generation to synthesize a training dataset for fine-tuning a smaller model (Ye et al., 2022a, b; Yehudai et al., 2024a). We refer to this type of dataset synthesis as FewGen.
FewGen relies on a generalist LLM’s ability to perform task inversion. For example, in Figure 1, we illustrate the task of detecting politically-biased articles. By using the prompt “Write a news article using harsh political language, a mocking tone and toxic commentary” and a few in-context learning examples, sampling continuations from an LLM is often sufficient to generate plausible news articles in the biased style we seek to detect. However, when thousands of completions are sampled from this prompt to construct a large dataset for distilling a specialist model, we observe repetition, bias towards popular entities, and stylistic differences from human-written texts. If the data is low diversity, a trained specialist model may not learn the task well.
In this work, we seek to alleviate the lack of diversity in synthetic data. We make the observation that dataset synthesis may be decomposed as two distinct LLM competencies: content sourcing, where the LLM obtains relevant information for the task, and task inversion, where the LLM generates a synthetic input example. Prior work has focused mainly on task inversion, i.e., generating good covariates using a target-conditioned prompt, while implicitly using the LLM’s parametric memory for content sourcing. In contrast, we investigate the importance of an explicit content sourcing stage.
We propose Synthesize by Retrieval and Refinement (SynthesizRR [^1] Pronounced “synthesizer”.), a procedure by which a large corpus acts as the guiding signal for the content sourcing stage. In the initial step, we repurpose each in-context learning input as a retrieval query. We use this to retrieve several documents from a large domain-specific corpus. We then use a generalist LLM for task inversion on each retrieved document. As a unique retrieved document is used in each prompt, the synthesis procedure yields more diverse examples, containing a wider range of real-world entities and assertions.
We benchmark SynthesizRR against FewGen on six text classification tasks, selected carefully to measure a variety of different styles of dataset synthesis. Experiments on an in-domain corpus (§5) reveal that SynthesizRR significantly surpasses FewGen in diversity and resemblance to human-authored texts, even though both procedures utilize the same frozen LLM. In §6, we see that student classifiers fine-tuned on SynthesizRR-generated data perform better than those fine-tuned on FewGen. We analyze the contributing factors in the student’s training dynamics, finding that our approach generates examples with superior class-separability, particularly for multiclass tasks.
Finally, in §7, we compare SynthesizRR to six prior approaches for synthesis of classification datasets: AttrPrompt Yu et al. (2023a), ReGen Yu et al. (2023b), SunGen Gao et al. (2023), ZeroGen Ye et al. (2022a), ProGen Ye et al. (2022b), and S3 Wang et al. (2023a). In general, we find that SynthesizRR gives higher diversity data, and in most cases, leads to higher student accuracy as a result.
Our contributions are as follows: (1) we propose a new method for integrating retrieval into example synthesis for teacher-student distillation; (2) we perform a detailed analysis comparing our method with existing approaches across both intrinsic data diversity metrics and downstream task accuracy.
2 Background and Task setup
The dataset generation tasks we focus on in this paper are in the domain of text classification. Denote an example as consisting of an input $x$, typically a natural language string, and output $y\in\mathcal{Y}$ for an output space $\mathcal{Y}$ of $C$ classes.
We assume access to a seed set of examples $\mathcal{D}_{\textsc{Seed}}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ which is large enough to cover the classes but small enough that a user could feasibly gather or construct these examples manually. In §5, we use the in-context learning set as $\mathcal{D}_{\textsc{Seed}}$.
Our assumption is that $\mathcal{D}_{\textsc{Seed}}$ is not large enough to train an effective student, and instead a larger synthetic dataset $\mathbin{\mathbin{\mathcal{D}_{\textsc{Synth}}}=\left\{(\mathbin{\tilde{x}^{i}},\mathbin{y^{i}})\right\}^{m}_{i=1}}$ ($m>>n$) will be needed. We repeatedly prompt a teacher language model $\mathcal{M}_{\text{LM}}$ to generate synthetic covariates $\tilde{x}$ given corresponding labels $y$ to produce $\mathbin{\mathcal{D}_{\textsc{Synth}}}$, a process we refer to as task inversion. A student language model $\mathcal{M}_{\text{S}}$ (e.g., a BERT-style pre-trained encoder model (Devlin et al., 2019)) is trained on $\mathbin{\mathcal{D}_{\textsc{Synth}}}$. We denote the student’s task (predicting $y$ from $x$) by $\tau$ and the teacher’s task (generating $x$ given $y$) by $\tau_{inv}$.
SynthesizRR aims to address the lack of diversity by leveraging retrieval from a large corpus $\mathcal{R}$ for the content sourcing step. Figure 2 illustrates how our method can be used to generate appropriate data. First, we start by executing a retrieval step based on the examples $\mathcal{D}_{\textsc{Seed}}$. Our assumption here is that we can find similar examples in the corpus, which are domain-similar enough to $\mathcal{D}_{\textsc{Seed}}$ to be usable as training examples. A context refinement prompt is then constructed to perform task inversion on each retrieved document. The LLM’s context window thus sees a unique and grounded prompt for each example it generates. This overcomes the need to memorize a corpus worth of information in the limited parameters of the teacher LLM.
We assume the existence of corpus $\mathcal{R}$, with each document containing information relevant to the task. However, the documents do not need to be unlabeled examples for our task; they can be more distantly related and still provide useful information for generating examples. In §5 we successfully generate customer reviews and questions about products, while retrieving from a corpus where documents are product titles and descriptions. Appendix C looks into how the corpus influences synthesis. Another reason raw retrieval documents are unsuitable for direct use as examples is due to the chunking process in retrieval, which can introduce coreference failures, among other problems (Mialon et al., 2023)).
Due to the difficulty of the task inversion step as a result of this mismatch between retrieved documents and test examples, we limit our investigation to chat-tuned teacher LLMs which have been shown to demonstrate strong instruction-following capabilities (Ouyang et al., 2022; Touvron et al., 2023; Bai et al., 2022).
Algorithm 1 SynthesizRR Retr-ICL
Input A set of seed examples $\mathcal{D}_{\textsc{Seed}}$, retrieval corpus =$\mathcal{R}=\{r_{k}\}$, retrieval model $\mathcal{M}_{\mathrm{ret}}$, expansion factor $K$, cosine-similarity criterion $(s_{\alpha},s_{\beta})$, teacher model $\mathcal{M}_{\text{LM}}$, prompt template $\mathcal{P}_{\tau}$, context refinement instruction $R_{inv}$, verbalizer $\mathcal{V}:\{y_{1},\dots,y_{C}\}\rightarrow\{v_{1},\dots,v_{C}\}$.
Output Synthetic dataset $\mathbin{\mathcal{D}_{\textsc{Synth}}}{}$
Procedure $\textsc{SynthesizRR}(\mathcal{D}_{\textsc{Seed}},\mathcal{R})$:
$\mathcal{D}_{\textsc{Retr}}\leftarrow\emptyset$
$\mathcal{D}_{\textsc{ICL}}\leftarrow\emptyset$
$\mathbin{\mathcal{D}_{\textsc{Synth}}}{}\leftarrow\emptyset$
$\triangleright$ Content sourcing using retrieval:
for $(x,y)\in\mathcal{D}_{\textsc{Seed}}$ do
     $[r_{1},\ldots,r_{K}]\leftarrow\mathcal{M}_{\mathrm{ret}}(x)$
     $\mathcal{D}_{\textsc{Retr}}\leftarrow\mathcal{D}_{\textsc{Retr}}\cup\{(x,y,[r_{1},\ldots,r_{K}])\}$
$\triangleright$ In-context learning set construction:
for $(x,y,[r_{1},\ldots,r_{K}])\in\mathcal{D}_{\textsc{Retr}}$ do
     for $r_{k}\in[r_{1},\ldots,r_{K}]$ do
         $\mathcal{D}_{\textsc{ICL}}\leftarrow\mathcal{D}_{\textsc{ICL}}\cup\{(r_{k},x)\}$ if $s_{\alpha}\leq\cos(x,r_{k})\leq s_{\beta}$
$\triangleright$ Task inversion:
for $(x,y,[r_{1},\ldots,r_{K}])\in\mathcal{D}_{\textsc{Retr}}$ do
     for $r_{k}\in[r_{1},\ldots,r_{K}]$ do
[…]
              $\mathbin{\tilde{x}}^{i}_{j}{}{\sim}{}\mathcal{M}_{\text{LM}}\left(\cdot|\mathbin{\tilde{x}}^{i}_{<j},\mathcal{P}_{\tau}(R_{inv},r_{k},\mathcal{V}(y)),\mathcal{D}_{\textsc{shots}}\right)$
         $\mathbin{\mathcal{D}_{\textsc{Synth}}}{}\leftarrow\mathbin{\mathcal{D}_{\textsc{Synth}}}{}\cup\{(\tilde{x}^{i},y)\}$
return $\mathbin{\mathcal{D}_{\textsc{Synth}}}{}$
3 Method
Algorithm 1 shows our dataset generation method.
Step 1. Retrieval using $\mathcal{M}_{\mathrm{ret}}$: SynthesizRR uses each in-context demonstration $(x_{{}_{\text{ICL}}},y_{{}_{\text{ICL}}})$ as a query for information retrieval, in addition to its role as an in-context demonstration for task inversion as discussed subsequently. When viewed as queries, these demonstrations lead to an expansion as we retrieve $K$ documents $\{r_{1},\dots,r_{K}\}$ of decreasing similarity for each in-context input $x_{{}_{\text{ICL}}}$. We rank contexts using retrieval model $\mathcal{M}_{\mathrm{ret}}$ and retain contexts with cosine similarity in (0.4, 0.9), where the upper limit helps avoid duplicates by excluding retrieved contexts if they are too close to the ICL example. We thus append each triplet $(x_{{}_{\text{ICL}}},y_{{}_{\text{ICL}}},[r_{1},\dots,r_{K}])$ to the set $\mathcal{D}_{\textsc{Retr}}$.
Step 2. Retrieval-augmented task inversion: The retrieved documents have lexical and semantic overlap with the query, but are not necessarily suitable examples for the final dataset. To accommodate this, we use a prompt to perform task-inversion using content sourced from each retrieved document, which we refer to as contextual refinement.
The minimum elements of the task inversion prompt $\mathcal{P_{\tau}}$ are the instruction $\mathcal{I}_{inv}$ and target $y$. We use a verbalizer function (Schick and Schütze, 2021; van de Kar et al., 2022) that provides a unique and descriptive representation of each label, i.e. $\mathcal{V}:\{y_{1},\dots,y_{C}\}\rightarrow\{v_{1},\dots,v_{C}\}$. We follow prior work on classification-based task inversion (Schick and Schütze, 2021; Ye et al., 2022a, b; Yu et al., 2023b; Gao et al., 2023) and use descriptive verbalizations to induce label-separability in the final dataset.
FewGen uses the standard causal language modeling objective to induce next-token probabilities from $\mathcal{M}_{\text{LM}}$. Nucleus sampling Holtzman et al. (2019) is used to autoregressively sample next tokens until the <eos> token is generated. This becomes synthetic example $\mathbin{\tilde{x}}^{i}$.
[…]
For each label $y$, we sample $m/C$ times to generate $\mathcal{D}_{\textsc{Synth}}=\bigcup\limits_{y\in\mathcal{Y}}{}\bigcup\limits_{i\in\{1..m/C\}}\left\{(\mathbin{\tilde{x}^{i}},y)\right\}$.
For SynthesizRR, the task inversion prompt is at minimum composed of the contextual refinement instruction $\mathcal{R}_{inv}$ alongside the verbalized target from the query example $\mathcal{V}(y_{{}_{ICL}})$. The LLM’s context window thus sees a unique and grounded prompt when auto-regressively generating each synthetic input $\mathbin{\tilde{x}}^{i}$:
[…]
for all documents $\text{for all documents }r_{k}\in\{r_{1},\cdots r_{K}\}$
We continue to use nucleus sampling to get diverse generations. Each in-context example thus produces $K$ unique synthetic examples $\{\tilde{x}_{1},\dots,\tilde{x}_{K}\}$, where we refer to $K$ as the “expansion factor”.
Step 3. Student distillation: The student is fine-tuned on $\mathbin{\mathcal{D}_{\textsc{Synth}}}$ by passing the BERT [CLS] token embedding through a feedforward layer. This produces a probability distribution over the labelspace $C$. We optimize the cross-entropy loss of the true label $y$ (which was used in the task-inversion prompt, thus is known to us) when the student model is passed the generated example $\mathbin{\tilde{x}}$. As we derive $\mathbin{\tilde{x}}$ from a teacher LLM, this can be considered a form of symbolic knowledge distillation (West et al., 2022) without the need for a critic model.
Unlike prior work, we make no changes to the student model to aid the out-of-distribution performance and focus solely on improvements to the dataset synthesis procedure.
In-context learning in SynthesizRR. Contextual refinement may also benefit from in-context demonstrations. In this case, we face a challenge in constructing an in-context learning set which accurately captures our context refinement task $r_{k}^{i}\rightarrow\mathbin{\tilde{x}}^{i}$. A basic exploration using handpicked demonstrations reveals that in-context learning primarily influences the format, rather than content, of synthetic examples. This finding is also supported by the literature on in-context learning (Min et al., 2022). We thus develop two approaches to in-context learning (ICL).
1. Retr-ICL: we use retrieval to construct a set of ICL examples $\mathcal{D}_{\textsc{ICL}}$, such that each ICL example looks similar to the SynthesizRR task-inversion prompt. We consider top-1 and top-2 retrieved results from dense retriever $\mathcal{M}_{\mathrm{ret}}$, and use a cosine-similarity criterion $s_{\alpha}\leq\cos(x_{{}_{\text{ICL}}},r_{k})\leq s_{\beta}$ to determine whether there is a possible match between the retrieved document $r_{k}$ and the ICL covariate $x_{{}_{\text{ICL}}}$. Within a massive corpus, there are typically sufficient semantically-close retrieved results to form an in-context pair ($r_{k},x_{{}_{\text{ICL}}}$). When used with an appropriate prompt template from Appendix E these ICL examples demonstrate the required format and approximately match in content.
2. No-Retr-ICL: a baseline method, where we use retrieval for content sourcing, but not for in-context learning. In No-Retr-ICL, for each generation we select 32 ICL examples at random from $\mathcal{D}_{\textsc{Seed}}$. Each ICL example is preceded by a prefix such as “News Article:” or “Product details:”, but without any instruction. After these ICL examples, we append the retrieved document $r_{k}$ and context refinement instruction $R_{inv}$ to produce the final prompt. The resulting in-context learning prompt is very similar to that used by FewGen, but also uses content-sourcing by including $r_{k}$ and $R_{inv}$. This baseline highlights the importance of using $\mathcal{D}_{\textsc{ICL}}$ in the Retr-ICL approach.
Dataset Class Train, Test In-domain corpus Task inversion
AG News 4 $115\text{k},7.6\text{k}$ RN/Dom Easy
Hyperpartisan 2 $516,65$ RN/Dom Medium
ToI Headlines 10 $52\text{k},10\text{k}$ RN/Ind Easy
Category 23 $30\text{k}^{*},2.4\text{k}$ Products Medium
Humor 2 $15\text{k},3\text{k}$ Products Hard
Polarity 2 $72\text{k}^{*},7.2\text{k}^{*}$ Products Medium
Table 1: Dataset statistics and our estimate of task inversion complexity. Downsampled for convenience.
4 Experimental Setup
Tasks and complexity. We experiment on 6 datasets described in Table 1, which are selected to give diverse task inversion steps. Previous work primarily benchmarks sentiment and topic classification datasets like AG News (Zhang et al., 2015), NYT (Meng et al., 2019), and Reddit (Geigle et al., 2021). We broaden from topic classification, which primarily involves summarization during the task inversion step, which LLMs are adept at (Goyal et al., 2022). Hyperpartisan (Kiesel et al., 2019) detects bias in political news, so the task inversion step includes a more substantial rewriting of neutral retrieved articles to form biased examples. Task inversion for Humor (Ziser et al., 2020), Category Blitzer et al. (2007) and Polarity (Ni et al., 2019) involves generating humorous questions and reviews from retrieved product details, which require additional skills.
Corpus Domain Docs Item Token
[…]
Table 2: Corpus statistics. Used LLaMa2 tokenizer.
Each task in Table 1 has at least one corpus which can be considered in-domain for the task. We ground the synthesis to this in-domain corpus during SynthesizRR. Prompts for all tasks are in Appendix E.
In §7, we compare to six prior approaches on four popular tasks: Yelp Polarity (Zhang et al., 2015), IMDb Maas et al. (2011), SST-2 Socher et al. (2013) and AG News, when grounding to smaller corpora (42k-150k documents). We view these tasks as less aligned with the goals our paper, so we do not focus our evaluation on them here.
Method Example
Gold There is decent bass, but the highs are a bit soft. A quick tweak to my equalizer, and they’re great. After reading several of the reviews on Amazon, I was a bit worried about the sound, but now that I have them I’m very happy. They’re a good price, and sooooo much better than the little ipod-like earbuds I’ve tried before. Those never stayed in my ear, and the bass never made me happy
FewGen I’ve been a very happy customer of this company for a long time. It is fast and does everything I need it to. I would definitely recommend it to anyone looking for a good external drive. However, I do have one issue with the product. The instructions that come with it are not very clear and I had a hard time figuring out how to properly use it.
(Retrieved Product) Portable Laptop Microphone. Connects to 1/8" mini microphone input on laptop. Right-angle shaped. Flat-frequency response.
SynthesizRR The portable laptop microphone is right-angled and has a flat-frequency response, making it easy to use for online meetings and interviews. It connects to the 1/8" mini microphone input on my laptop and has worked great for the past two months, but I have noticed some distortion in the audio when I move around too much. Overall, it’s a great value for the price and has made my remote work and video conferencing much more productive and efficient.
Table 3: Real and synthetic examples from “electronics” class of Category. Grey text indicates lack of specifics.
Retrieval corpora. Table 2 describes corpora used for retrieval. We consider four corpora in different domains, each with millions of records: RealNews/Dominant (US/EU News), RealNews/Regional (Regional News), RealNews/India (Indian News), Products (Amazon products metadata (Ni et al., 2019)).
Retrieval model. We use Contriever (Izacard et al., 2022) for dense retrieval from each corpus. This performs a semantic match between the query and each document using cosine-similarity. In Appendix B, we also perform an ablation study using sparse retriever BM25, which does lexical matching between each query-document pair.
Teacher LLMs. As teacher models, we primarily use a frozen LLaMa-2 Chat 13B (Touvron et al., 2023) for the task inversion step in SynthesizRR and FewGen. We also experiment with ClaudeV1 (Appendix H). For in-context learning (ICL) Brown et al. (2020), we select examples randomly from the train set: 50 ICL examples/class for multi-class and 100/class for binary tasks. We believe this is a realistic number of examples that a system designer could source if they were to put some effort into building a specialist model. We pick these numbers to provide sufficient ICL diversity in 32-shot FewGen and SynthesizRR.
Student model. Specialization performance is measured on student LM DeBERTa-v3-Large (435M params, He et al. (2021)).
Figure 3: Self-BLEU ()($\downarrow$)( ↓ ) for ngrams n=1-5. Comparison: Gold, FewGen 0-shot, FewGen 32-shot, SynthesizRR 0-shot, SynthesizRR 3-shot Retr-ICL, SynthesizRR 32-shot No-Retr-ICL.
Evaluation criteria. Text generation can be challenging to evaluate objectively in multi-task scenarios (Chang et al., 2024). Therefore in §5 we evaluate synthetic text based on several criterion, to detect behaviours we observe during synthesis as in Table 3. Self-BLEU (Papineni et al., 2002; Zhu et al., 2018) measures lexical diversity of the dataset based on $n$-gram overlap between pairs of examples. Entity entropy measures the diversity of entities using the probability distribution of each of 16 entity-types (inferred from a pretrained named entity recognition model). Dataset which over-represent popular entities score lower on entropy. On the other hand, Entity recall and Entity KL divergence compares the similarity of entities compared to Gold, and datasets which reproduce entities frequently seen in Gold data score higher. MAUVE (Liu et al., 2021) measure similarity to human-written text using pretrained representations from a GPT2-XL model; this indicates distributional differences that would be perceived by language models.
Besides the final accuracy of the student model, we also consider label preservation accuracy, which is obtained from an “oracle” model for the task. We construct the oracle from Gold data by running a grid-search over DeBERTa-v3-Large hyperparams, splitting 80% of the Gold train set for fine-tuning and 20% for validation. Then, we measure the fraction of synthetic examples which the oracle classifies to belong to the prompted target class. This indicates the adherence of the generated example to the class it should belong to, as per the prompt.
Figure 4: Entity entropy ()($\uparrow$)( ↑ ) on ToI (headlines) and Category (reviews). Comparison: Gold, FewGen 32-shot, SynthesizRR 3-shot Retr-ICL and SynthesizRR 32-shot No-Retr-ICL. Zero-shot results are similar for SynthesizRR and worse for FewGen; we omit them.
5 Results: Intrinsic Evaluation
In this section we focus on evaluating intrinsic properties of the generated datasets, including their diversity and entity coverage. We focus on a LLaMa-2 Chat 13B teacher LLM. We generate datasets of a fixed size comparable to the number of Gold rows: $8$k rows (AG News, ToI Headlines, Category), $4$k rows (Polarity) or $2$k rows (Hyperpartisan (516 originally), Humor). Examples of generations on each dataset can be found in Appendix F.
RQ: Does retrieval augmentation improve lexical diversity? Figure 3 shows lexical diversity within the dataset. Human-written texts (Gold) score high on lexical diversity (low Self-BLEU). FewGen texts tend to reuse the same words and phrases, leading to repeated text across generations (high Self-BLEU). SynthesizRR text has lexical diversity approaching human text for all n-gram values. We note in-context learning has an inconsistent effect; it improves the lexical diversity for news corpora but not for products.
RQ: Does SynthesizRR address entity diversity? Popularity bias is a phenomenon wherein LLM generations tend to over-represent popular “head” entities. It has been studied for QA tasks (Mallen et al., 2023; Kandpal et al., 2023).
In Figure 4 we see how SynthesizRR eliminates popularity bias in several entity types at once. Here, we use Spacy NER model en_core_web_lg (Honnibal et al., 2020) to detect entity occurrences in each synthetic example. By sourcing from the long-tail of retrieval results ($k=50$), the generated dataset has much higher entity entropy, meaning the probability distribution of each entity-type is closer to a uniform distribution compared to FewGen. This positions SynthesizRR closer to Gold, which also shows high entity entropy.
Method Norp Org Person Gpe Recall ()($\uparrow$)( ↑ ) KL div. ()($\downarrow$)( ↓ )
Unique Entities
Gold 319 3943 3952 712 - -
FewGen* 43 480 400 73 0.05 -
SynthzRR 137 2718 1528 238 0.12 -
SynthzRR 109 1755 1012 178 0.10 -
Total Entities
Gold 843 7233 6096 1558 - -
FewGen* 94 775 506 96 0.23 3.10
SynthzRR 319 3991 1989 397 0.35 2.35
SynthzRR 314 2699 1464 363 0.32 2.52
Table 4: Entity similarity in Category ($8$K). We show the counts of unique and total entities for 4 entity-types. Entity recall measures the fraction of Gold entities co-occuring in the synthetic data; in the bottom half, we additionally weigh each entity by its frequency in Gold. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
RQ: How is entity similarity in synthetic data affected by grounding to an in-domain corpus? Generating diverse entities is not beneficial if they are not also present in Gold. We thus study entity similarity. For the Category task we generate $8$k product reviews and randomly select $8$k Gold examples. In Table 4, we measure entity recall, and find that the occurrence of Gold entities is 100%-140% higher in SynthesizRR than FewGen. The KL divergence of each entity-distribution is also lower. We finally consider the entity coverage (unique entities) and entity density (total entities). Compared to Gold, FewGen tends to produce fewer unique entities (places, events, languages, currencies, etc). Each FewGen example also has a lower density of entities, as visible in Table 3. SynthesizRR coverage and density more closely match Gold.
RQ: How distributionally similar are our generated examples and human-written examples? We calculate MAUVE score, which uses GPT-2XL (Radford et al., 2019) to embed texts and compare them based on Type-1 and Type-2 errors (Liu et al., 2021). We see from Table 5 that zero-shot generations are quite dissimilar in both approaches compared to few-shot methods.
Method AG. Hyp. ToI Cat. Hum. Pol.
(Dataset size) ($8$K) ($4$K) ($8$K) ($8$K) ($2$K) ($4$K)
Zero shot
FewGen $56.57$ $53.68$ $62.79$ 63.2 $75.6$ $62.82$
SynthzRR 90.3 59.2 63.0 $61.06$ 82.9 78.6
Few shot
FewGen* $56.7$ $65.39$ $60.33$ $65.8$ $78.14$ $69.21$
SynthzRR 92.0 72.8 87.9 75.2 87.5 89.9
SynthzRR $91.76$ $67.86$ $67.16$ $75.14$ $86.95$ $83.2$
Table 5: MAUVE similarity score ($\uparrow$) using GPT2-XL embeddings. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
Method Teacher LM AG. Hyper. ToI Categ. Humor Polar. Avg
(Dataset size) ($8$K) ($4$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold 91.015 78991.01578991.015\,78991.015 789 93.230 76993.23076993.230\,76993.230 769 $82.488$ 81.516 66781.51666781.516\,66781.516 667 $93.06$ $95.2645$ 89.43
Zero shot
FewGen LLaMa2 69.457 89569.45789569.457\,89569.457 895 72.6 $32.09$ $62.35$ 74.433 33374.43333374.433\,33374.433 333 $80.996$ 65.32
SynthesizRR LLaMa2 83.510 52683.51052683.510\,52683.510 526 69.846 15469.84615469.846\,15469.846 154 74.4 68.9 82.5 $84.707$ 77.32
FewGen ClaudeV1 75.031 57975.03157975.031\,57975.031 579 57.538 46257.53846257.538\,46257.538 462 $23.298$ $47.1$ 49.866 66749.86666749.866\,66749.866 667 $87.481$ 56.72
SynthesizRR ClaudeV1 83.9 72.307 69272.30769272.307\,69272.307 692 $71.83$ 66.783 33366.78333366.783\,33366.783 333 62.113 33362.11333362.113\,33362.113 333 88.7 74.29
Few shot
FewGen* LLaMa2 $84.2$ 74.461 53874.46153874.461\,53874.461 538 73.7 68.641 66768.64166768.641\,66768.641 667 $88.38$ $90.899$ 80.05
SynthesizRR LLaMa2 82.992 10582.99210582.992\,10582.992 105 78.461 53878.46153878.461\,53878.461 538 $73.262$ 72.4 90.2 $91.0015$ 81.38
SynthesizRR LLaMa2 85.2 79.1 $72.826$ 71.941 66771.94166771.941\,66771.941 667 88.773 33388.77333388.773\,33388.773 333 $88.197$ 81.00
FewGen* ClaudeV1 75.852 63275.85263275.852\,63275.852 632 58.461 53858.46153858.461\,53858.461 538 $72.224$ 68.833 33368.83333368.833\,33368.833 333 $82.94$ $91.2435$ 74.93
SynthesizRR ClaudeV1 $83.7394$ 72.307 69272.30769272.307\,69272.307 692 $72.832$ 65.408 33365.40833365.408\,33365.408 333 83.393 33383.39333383.393\,33383.393 333 91.3 78.16
SynthesizRR ClaudeV1 83.710 52683.71052683.710\,52683.710 526 $72.00$ $72.45$ 67.816 66767.81666767.816\,66767.816 667 $76.2066$ $87.87$ 76.68
Table 6: Test Accuracy ($\uparrow$) after distilling DeBERTa-v3-Large student from LLaMa-2 Chat 13B and Claude Instant-v1. We report mean of 5 runs and rerun all 5 in cases of std $\geq$ 6.0. The top half considers zero-shot synthesis and bottom half uses in-context learning, where we bold the best result under each paradigm. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
Surprisingly, SynthesizRR generations are much more similar to human text than FewGen. This is surprising as nothing in our content sourcing strategy explicitly guides SynthesizRR generations to match the distribution of Gold. We thus manually inspect generations and discover an interesting pattern which can be attributed to content sourcing. As shown earlier, and in Table 3, the density of entities is higher under SynthesizRR. FewGen produces generations which obey the prompt, but are very bland and do not include specifics. On the other hand, by obtaining information-rich documents, SynthesizRR is able to ground the task inversion step in details of the retrieved article/product. We hypothesise that this improves the MAUVE score towards Gold, which is similarly grounded in specifics.
We further analyze the effect of the retrieval corpus in Appendix C.
6 Results: Student distillation
We have established that SynthesizRR generates more diverse datasets compared to a baseline approach. Now, we return to the application of training a specialist model based on these datasets.
Table 6 shows the results of training a DeBERTa-v3L student on generated datasets across several settings. We report the mean of 5 runs and rerun if the standard deviation is $\geq$6(indicating one or more models failed to train). In the zero-shot setting, we find that SynthesizRR performs much better than FewGen, despite using the same frozen teacher LLM. Note that SynthesizRR uses in-context examples for retrieval here whereas FewGen does not; our method has some additional supervision here. However, in this setting, we see clear gains during the task inversion stage ($\uparrow$12for LLaMa2 and $\uparrow$17.6for ClaudeV1). Thus, having access to retrieval yields a better final dataset, almost on par with the few-shot generation.
With in-context learning, 3-shot SynthesizRR using the No-Retr-ICL strategy trains better students than 32-shot FewGen ($\uparrow$1.3for LLaMa2 and $\uparrow$3.2for ClaudeV1). We conclude that naively adding ICL examples is not an effective use of the LLM’s context window. Instead, a better content sourcing strategy positively influences the student distillation, leading to faster synthesis and better test performance.
Analysis of student distillation performance. Why does SynthesizRR improve classification dataset synthesis? We take a closer look at the few-shot generation strategies from a student training perspective.
We first look at the label preservation accuracy, obtained from an oracle DeBERTa-v3L model tuned on the full Train dataset (Appendix H). We would expect that better label preservation means a higher-fidelity training dataset. However, Table 7 shows that FewGen datasets have very high label preservation in spite of their lower test performance. Especially on multiclass tasks (AG., ToI, Cat.), FewGen shows the highest label preservation (exceeding Gold) but this does not translate into improved student performance.
Figure 5: Data maps from a DistilBERT training run on $8$K Category rows from LLaMa2. FewGen (center) is skewed towards easy-to-learn examples (top-left) while Gold (left) and SynthesizRR (right) have a higher density of ambiguous examples.
To understand this, we conduct a deeper analysis of the student training dynamics on multiclass datasets. We train a DistilBERT student for 6 epochs and plot the corresponding data-maps Swayamdipta et al. (2020). As noted in §4, label preservation can be exploited by only generating easy to classify examples. From the Category data maps in Figure 5, it is clear that FewGen generations tend to cluster around easy-to-learn examples (high confidence and low variability). SynthesizRR contains more ambiguous examples (high variability) which we hypothesize are essential to learning the nuances between classes. To validate this, we filter out the least ambiguous examples (bottom 17% variability) and retrain the students. In Table 8 we find that FewGen performance degrades, whereas synthesizer improves (giving us new best performances, despite using only 83% of rows). We conclude that SynthesizRR generates more ambiguous examples, and this helps establish better class-separability in multi-class data sets.
Method AG. Hyp. ToI Cat. Hum. Pol.
(Dataset size) ($8$K) ($4$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold $93.81$ $81.55$ $85.23$ $84.8$ $95.5$ $96.6$
LLaMa2 Few shot
FewGen* 92.4 $71.25$ 85.9 88.1 $71.7$ $94.75$
SynthzRR $86.88$ 78.6 $74.29$ $72.11$ $90.65$ $94.77$
SynthzRR $87.64$ $75.5$ $74.85$ $74.46$ 95.7 97.6
ClaudeV1 Few shot
FewGen* 94.5 $63.75$ 87.4 89.4 $85.85$ $99.55$
SynthzRR $87.56$ 72.8 $74.79$ $69.41$ 90.7 $99.32$
SynthzRR $87.36$ $65.85$ $73.24$ $73.24$ $77.35$ 99.7
Table 7: Few-shot label-preservation accuracy ($\uparrow$) using tuned oracle DeBERTa-v3L model. Gold row is accuracy on 20% validation split. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
Method AG. ToI Cat. Avg
(Dataset size) ($6.6$K) ($6.6$K) ($6.6$K)
LLaMa2 Few shot
[…]
ClaudeV1 Few shot
[…]
Table 8: Test Accuracy ()($\uparrow$)( ↑ ) after keeping 83% most-ambiguous examples. We report improvements compared to Table 6. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
7 Comparison to additional baselines
Method Teacher / Retriever Data Size Yelp AG. IMDb SST-2
Gold - 6k 94.5 90.8 91.3 88.2
ZeroGen GPT2-XL 6k
$\otimes$
$\otimes$
[…]
82.4
(Reported) 200k
$\otimes$
$\otimes$
84.3 87.3
ProGen GPT2-XL 6k
[…]
(Reported) 100k 89.4 80.8 84.1 87.2
SunGen GPT2-XL 6k
[…]
84.9
[…]
(Reported) 200k 89.2 80.5 89.4 89.5
ReGen BERT 6k
[…]
82.7
[…]
(Reported) 20k 89.0 85.0 87.8 85.3
S3 GPT3.5-T 6k
$\otimes$
$\otimes$
87.1
$\otimes$
(Reported) 21.2k
$\otimes$
$\otimes$
89.0
$\otimes$
AttrPmt GPT3.5-T 6k 89.3 79.8
[…]
80.8
(Reported) 6k 90.6 82.7 85.9 83.6
Zero shot
SynthzRR LLaMa2 $+$ Contr. 6k 85.6 85.3 82.9 80.2
SynthzRR ClaudeV1 $+$ Contr. 6k 87.5 85.6 83.6 82.5
SynthzRR LLaMa2 $+$ BM25 6k 83.2 84.3 74.1 84.4
SynthzRR ClaudeV1 $+$ BM25 6k 86.7 84.2 79.1 82.6
3-shot Retr-ICL
SynthzRR LLaMa2 $+$ Contr. 6k 89.9 84.6 84.8 83.8
SynthzRR ClaudeV1 $+$ Contr. 6k 91.3 86.0 86.3 80.6
SynthzRR LLaMa2 $+$ BM25 6k 90.7 84.3 84.7 84.4
SynthzRR ClaudeV1 $+$ BM25 6k 89.4 84.1 81.6 82.3
Table 9: Test accuracy of SynthesizRR Retr-ICL against baselines after fine-tuning a DistilBERT student. For each released dataset, we subsample to 6k examples to make a fair comparison. We train 5 student models and report the mean accuracy (std. dev. was $\leq 2.0$ in all cases, thus we avoid reporting it for brevity). We report the DistilBERT accuracy from the original paper (if available) in italics. For each dataset, we bold the best result from the fair evaluation. Tasks not evaluated by previous authors are denoted by
$\otimes$
, while those evaluated without dataset release are marked as
[…]
. GPT3.5-T is gpt-3.5-turbo.
Here, we explore how SynthesizRR directly compares to prior work on synthesis of text classification datasets. We compare against six prior approaches:
ZeroGen Ye et al. (2022a): a simple approach which uses zero-shot generation from GPT2-XL under different prompting schemes. For certain tasks like IMDb and SST-2, the authors occasionally seed the prompt with a previously-generated movie name (described subsequently).
ProGen Ye et al. (2022b): generates a small synthetic dataset, then uses a student model along with an influence function to highlight high-quality generations. These are used as in-context examples for another round of generation.
SunGen Gao et al. (2023): uses the ZeroGen strategy to generate a large synthetic dataset (200k rows). Then, uses a custom bi-level optimization algorithm (involving the student model) to determine instance-weights of each synthetic example.
ReGen Yu et al. (2023b): performs multi-round filtering of retrieved results using 2 BERT models; one trained for retrieval, and one classifier. Use consistency between these models to filter noisy data.
S3 Wang et al. (2023a): Constructs a “seed dataset” (different from ours) and trains a student model. Then, extrapolate errors using an LLM and synthesizes additional data. We combine this with the seed data and repeat the process.
AttrPrompt Yu et al. (2023a): a method focused on improving diversity and unbiasedness of generated datasets. Prompts a powerful LLM like GPT3.5-Turbo with different attributes, each along different dimensions. Attributes are extracted from a human-in-the-loop analysis of task using GPT3.5-Turbo.
We note that zero shot generation and few-shot generation has already been compared in Table 6, so we do not include it here.
We benchmark four classification tasks which are popular in prior work: Yelp Polarity (Zhang et al., 2015), IMDb Maas et al. (2011), SST-2 Socher et al. (2013) and AG News (Zhang et al., 2015). The first three tasks are binary sentiment analysis, whereas the last is multi-class topic classification.
For the content sourcing stage of SynthesizRR, we retrieve documents from the following corpora:
  • Yelp Polarity: to generate positive and negative reviews about businesses, we retrieve from the Yelp Dataset corpus Yelp (2022) which contains details of 150k businesses. Each document contains the the name, address, and various tags and boolean attributes for a single business.
  • IMDb and SST-2: to generate movie reviews, we retrieve from the CMU Movie Summary corpus Bamman et al. (2013), which contains 42k plot summaries. Each document is a plot summary for a single movie.
  • AG News: we use RealNews/Dominant from Table 2.
We note that the CMU Movie Summary and Yelp Dataset corpora are between one to three orders of magnitude smaller than corpora in Table 2, thus it helps us evaluate SynthesizRR with smaller corpora. We include two different retrieval approaches: Contriever (semantic match) and BM25 (lexical match).
Following Yu et al. (2023a), we generate 6,000 rows using the SynthesizRR Retr-ICL variant described in Algorithm 1. We generate an equal number of examples for each class. We calculate intrinsic metrics Self-BLEU-5, MAUVE and Entity Entropy as described in §4.
Prior work often generates much larger datasets (20k to 200k examples) and uses different student model hyperparameters. Intrinsic evaluations of dataset quality are also seldom reported. This makes it difficult to fairly compare results from previous papers. Thus, we reproduce results ourselves by using synthetic datasets released by the authors, after subsampling to 6,000 examples. We report distillation results in Table 9 and intrinsic metrics in Table 10.
Method Teacher Retriever Self-BLEU-5 ()($\downarrow$)( ↓ ) Entity Entropy ()($\uparrow$)( ↑ ) Mauve ()($\uparrow$)( ↑ )
(Dataset) LM Yelp AG. IMDb SST-2 Yelp AG. IMDb SST-2 Yelp AG. IMDb SST-2
Gold - - 27.5 17.1 27.9 35.5 7.1 6.6 7.5 3.2 - - - -
ZeroGen GPT2-XL -
$\otimes$
$\otimes$
[…]
15.3
$\otimes$
$\otimes$
[…]
4.9
$\otimes$
$\otimes$
[…]
52.0
ProGen GPT2-XL -
[…]
SunGen GPT2-XL -
[…]
15.4
[…]
4.9
[…]
68.7
[…]
ReGen - BERT
[…]
56.5
[…]
8.1
[…]
68.1
[…]
S3 GPT3.5-T -
$\otimes$
$\otimes$
62.2
$\otimes$
$\otimes$
$\otimes$
5.7
$\otimes$
$\otimes$
$\otimes$
62.0
$\otimes$
AttrPmt GPT3.5-T - 70.3 39.8
[…]
71.5 3.4 6.0
[…]
3.4 55.9 52.8
[…]
50.0
Zero shot
SynthzRR LLaMa2 Contr. 70.2 29.3 66.3 41.9 5.4 7.1 5.7 4.5 67.9 89.5 58.5 50.0
SynthzRR ClaudeV1 Contr. 70.3 31.5 51.5 45.3 5.5 6.6 5.3 4.8 54.0 94.2 55.9 50.0
SynthzRR LLaMa2 BM25 70.6 28.7 62.2 36.5 4.7 7.0 5.6 5.1 64.7 90.3 60.5 50.0
SynthzRR ClaudeV1 BM25 72.9 30.9 50.4 36.9 4.5 6.5 5.1 5.4 54.6 90.8 53.2 50.0
3-shot Retr-ICL
SynthzRR LLaMa2 Contr. 59.5 34.2 62.9 26.3 4.9 7.2 5.7 3.8 84.9 92.6 72.6 50.0
SynthzRR ClaudeV1 Contr. 58.5 23.7 38.0 24.6 4.9 6.7 5.9 4.3 70.3 95.8 58.0 50.0
SynthzRR LLaMa2 BM25 66.6 32.0 59.7 25.3 5.1 7.2 5.6 4.8 81.9 92.5 78.7 50.0
SynthzRR ClaudeV1 BM25 59.9 24.6 41.9 26.8 5.0 6.7 5.4 4.9 63.3 96.0 58.5 50.0
Table 10: Intrinsic evaluations of synthetic datasets released by prior work. We subsample all to 6k examples before computing metrics as described in §4. Within each dataset, we bold the best result. Tasks not evaluated by previous authors are denoted by
$\otimes$
, while those evaluated without dataset release are marked as
[…]
. GPT3.5-T is gpt-3.5-turbo.
RQ: How does SynthesizRR compare to existing approaches in terms of distilled student accuracy?
We measure accuracy of a DistilBERT Sanh et al. (2019) student model. DistilBERT is a standard model used in prior dataset synthesis work Yu et al. (2023a); Ye et al. (2022a); Gao et al. (2023); Wang et al. (2023a); Ye et al. (2022b). We set the same model hyperparams as Yu et al. (2023a), i.e. training for 5 epochs using using Adam optimizer with batch_size=32 , lr=2e-5 , weight_decay=1e-4 and epsilon=1e-6 . We do not use cartography-based filtering for SynthesizRR to make a fair comparison. In addition to our results on 6k examples, we report original numbers from prior work (which might be on much larger datasets).
The results in Table 9 show that methods which rely on relatively weak LLM teachers like GPT2-XL Radford et al. (2019) can perform well on sentiment and topic tasks like IMDb, but require a very high data cost (15-30x more synthetic data than SynthesizRR). When scaled down to 6k rows, the performance of methods like ZeroGen and SunGen deteriorates.
Approaches which use strong instruction-following LLMs like AttrPrompt, S3, and SynthesizRR can achieve similar or better performance with much smaller datasets. These methods create high-quality datasets rather than modify the student modeling process, as is done by SunGen and ProGen. SunGen performs an iterative bi-level optimization over the ZeroGen datasets, jointly learning instance-weights and improving the student. ProGen uses an iterative filtering approach by training a student model and filtering partial generations based on a noise-robust, to curate an in-context dataset. We hypothesize these additions of the student model into the synthesis process also impact the final classification accuracy, as the dataset becomes specialized to the particular choice of student and its hyperparams. As we use a standard student-distillation setup here, datasets from these approaches may not perform as well.
More complex prompting techniques like Chain-of-Thought (Wei et al., 2022) used by S3 can indeed improve the task-inversion step, though this requires much higher API costs due to longer output lengths. Chain-of-Thought prompting thus seems like a promising approach by which SynthesizRR’s task-inversion step could be augmented.
RQ: do we find evidence that content sourcing promotes diversity and similarity?
In Table 10, we measure the entity diversity (Entity Entropy), lexical diversity (Self-BLEU), and similarity to Gold texts (MAUVE) compared to prior approaches. Among prior approaches, only AttrPrompt (Yu et al., 2023a, Appendix E) attempts to improve diversity of the generated text, by templatizing the task inversion instruction with attributes such as style, topic, length:min-words and more. ReGen is the only synthesis approach which uses content sourcing (but not task inversion). We thus consider these the two most relevant baselines for SynthesizRR.
We see that both ReGen and SynthesizRR achieve very high entity entropy compared to AttrPrompt, thus underscoring the importance of a content sourcing step. Unlike SynthesizRR, ReGen uses only retrieval, thus it has no explicit task-inversion step to make the contexts similar to Gold texts. Thus, we observe that ReGen suffers in terms of lexical diversity, Mauve and student accuracy, compared to SynthesizRR.
On the other hand, Chain-of-Thought prompting (S3), despite generating a strong classification dataset as we saw in Table 9, suffers from the lack of lexical diversity and similarity to Gold texts. This is also seen in AttrPrompt and previously in FewGen. This lends evidence to the hypothesis that synthesis without content sourcing tends to produce datasets with lower diversity, which cannot be overcome by complex prompting strategies alone.
Finally, we observe that ZeroGen shows very high diversity on SST-2, a task which involves generating partial movie reviews having positive or negative sentiment. However, as mentioned in Ye et al. (2022a, Section 4.6), ZeroGen is not simply zero-shot generation; the authors first generate a movie using the prompt Movie: and then insert the generated movie name into the prompt template The movie review in positive sentiment for movie "<Movie>" is: to generate an SST-2 example. We posit that the generated movie fulfils the same purpose as a retrieved context in SynthesizRR. Similarly high diversity is observed by SunGen which starts with ZeroGen-generated dataset and learns instance-weights (Gao et al., 2023, Section 2.2). Such synthetic content sourcing could be a fruitful direction for future work, to overcome the need for a retrieval corpus in SynthesizRR.
8 Related Work
Dataset synthesis using LLMs. Using LLMs to perform task inversion for dataset synthesis has been studied previously. Most use GPT-2XL without fine-tuning (Ye et al., 2022b, a; Gao et al., 2023; Meng et al., 2022; Schick and Schütze, 2021; Jung et al., 2023). Only recent works have considering large teacher LLMs such as GPT-3 (West et al., 2022; Honovich et al., 2023; Wang et al., 2023b) and chat-tuned LLMs such as gpt-3.5-turbo (Yu et al., 2023a; Yehudai et al., 2024b; Wang et al., 2023a).
For the generation of text classification datasets, class-conditioned prompting is key. Prior approaches have investigated zero-shot (Ye et al., 2022a), iterative few-shot prompting (Ye et al., 2022b), or synthesis using seq2seq LLMs fine-tuned on a curated dataset (Lee et al., 2021), but only recently has the importance of varying the prompt been established. Yu et al. (2023a) use prompts which explicitly attribute variations (such as subtopics and brands) and find this obtains greater diversity. We explore adding retrieval contexts to the prompt as the source of diversity and knowledge-grounding, and find it can significantly improve representation of tail entities. This avoids issues related to memorization of parametric knowledge (Roberts et al., 2020; Kandpal et al., 2023). We note that distillation for generation tasks like paraphrasing and summarization (Jung et al., 2023) has been demonstrated using off-the-shelf critic models; instead, our work focuses on classification and does not need a filtering step to be effective (though, as we show in §6, it does benefit from such steps in the multi-class case).
Retrieval-augmented generation. Our approach has many of the characteristics of in-context retrieval-augmented generation (RAG) (Lewis et al., 2020; Ram et al., 2023; Huang et al., 2023; Izacard et al., 2023). Previous studies show how RAG bypasses numerous problems associated with generating solely from parametric memory, i.e. heightened bias towards “head” entities (Mallen et al., 2023), lower lexical diversity (Holtzman et al., 2019; Jentzsch and Kersting, 2023), and hallucinated and confabulated information (Zhang et al., 2023). To our best knowledge, using retrieval-augmented generation for synthesis of classification tasks has not been explored at the instance level. Yu et al. (2023b) studied the retrieval-only setting for creation of topic and sentiment datasets; as described in §section 2 we consider tasks such as Hyperpartisan, Humor, and ToI Headlines, which pose additional challenges. Viswanathan et al. (2023) perform dataset-level retrieval and not instance-level retrieval, additionally selecting the student model based on the task whereas we focus on a fixed student. Finally, note that our application looks very different from other RAG applications like question answering (Asai et al., 2023) or counterfactual data augmentation (Paranjape et al., 2022); we generate diverse outputs only loosely connected to the input query.
Diversity in dataset synthesis. An important distinction between RAG and SynthesizRR is the desire for diverse generations. RAG has so far been used primarily for knowledge-based tasks such as open-domain question-answering, fact-checking and few-shot classification (Asai et al., 2023). These tasks face the issue of multiple valid surface-forms, which only change the wording of the generated output but not its meaning, making evaluation difficult (Holtzman et al., 2021). Greedy sampling is used most often.
On the other hand, dataset synthesis is emboldened by diverse surface-forms, as this produces diverse covariates linked to the same target, which intuitively aids generalization; we thus use top-p sampling (Holtzman et al., 2019). Another major difference from RAG is that we retain the long-tail of retrieved results, as this allows us to create a diverse grounded dataset.
9 Conclusion
In this work we describe how a retrieval corpus can be used to aid the synthesis of a text classification data set in specialized domains. We show that the diversity of the generated data is enhanced by including retrieved documents in a generation prompt. Compared to few-shot generation, we find that SynthesizRR produces more diverse and representative text and leads to better students.
Limitations
Most principally, our work relies on the existence of a large corpus that is close enough to the task at hand. This may be prohibitive for doing dataset generation in low-resource languages, where a large corpus of related content may not be available. It would be intriguing to explore cross-lingual transfer of content sourcing, but this would require additional experimental validation. By contrast, approaches like FewGen do not require this corpus.
The need for an explicit context sourcing step and increased prompt-length causes an increase in the expenses and latency, especially when using LLM APIs. Such increased expense may not be worth it in the presence of a poor quality retrieval corpus. For one, if the in-context examples are not easily reusable as queries, then SynthesizRR can retrieve irrelevant documents which might not be suitable for task inversion. Furthermore, in the case of factually dubious corpus documents, the student model may end up grounding in factually incorrect information. This can be mitigated by a human-in-the-loop step to remove such documents before task inversion.
Finally, we note that the scope of our experiments is restricted to a set of classification tasks over a few English domains of text. While we believe our approach can be applied to other languages, other domains, and tasks like question answering that go beyond classification, we have not validated this in this work.
Appendix A Risks
Although the main goal of our work is to improve text classification, our use of LLMs to generate examples does carry some conceptual risks. By generating news articles to train classifiers on, we run the risk of generating fake news and other harmful content. However, we believe this risk is mitigated by the fact that the final outcome of our system is a classifier: classification models have relatively constrained failure modes (misclassification) compared to text generation models that can mislead users. Furthermore, we do not believe our approach uniquely advances the generation of fake news; our advances are largely orthogonal to the technology that brings such risks.
Appendix B Dense vs sparse retrieval in SynthesizRR
Retriever AG. Hyp. ToI Cat. Hum. Pol. Avg.
(Size) ($8$K) ($4$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold 91.0 93.2 82.5 81.5 93.1 95.3 89.43
LLaMa2 Zero shot
Contr. 83.5 69.8 74.5 68.9 82.5 84.7 77.32
BM25 83.2 74.2 70.7 57.6 78.5 85.4 74.93
ClaudeV1 Zero shot
Contr. 83.9 72.3 71.8 66.8 62.1 88.7 74.29
BM25 83.2 57.2 69.8 53.7 73.9 91.8 71.60
LLaMa2 3-shot Retr-ICL
Contr. 83.0 78.5 73.3 72.4 90.2 91.0 81.38
BM25 82.1 77.9 71.9 65.4 87.5 87.4 78.69
ClaudeV1 3-shot Retr-ICL
Contr. 83.7 72.3 72.8 65.4 83.4 91.3 78.16
BM25 83.0 73.5 70.0 52.4 82.4 90.7 75.34
Table 11: Test accuracy after distilling a DeBERTa-v3L student on a dataset generated by SynthesizRR comparing BM25 and Contriever. We use the same seed set and corpus as Table 2. We train 5 student models and report the mean accuracy, rerunning all 5 in case of std $\geq$ 6.0. The top two subsections consider zero-shot synthesis and bottom two considers 3-shot Retr-ICL variant. We bold the best result in each subsection. Contriever numbers are reproduced from Table 6.
So far, a single dense retriever (Contriever) has been used for the content sourcing step by using a bi-encoder approach (Lee et al., 2019; Chen et al., 2017). We embed both the input in-context covariate and each corpus document, and then rank results based on cosine similarity. In §5, we retrieved $k=500$ documents for each in-context example and after filtering, randomly sampled among these to produce a grounded set of documents on which we apply our task inversion strategy Retr-ICL.
In this section we explore how changing the retrieval model affects the content sourcing stage and its downstream effects. Table 11 compares the DeBERTa-v3L accuracy of Contriever and BM25. Keeping other parts of the process the same, we switch Contriever to BM25 Okapi (Robertson and Zaragoza, 2009), a popular sparse retrieval method which performs a lexical match based on inverse term frequencies. Unlike dense retrieval models, BM25 outputs a score which is an unbounded positive number. Thus we are unable to use meaningful thresholds to bound the similarity in our Retr-ICL approach. Instead, we construct the Retr-ICL in-context set using the top-2 retrieved contexts for each ICL example and without applying the filter.
We observe in Table 11 that a powerful student model trained on the dense-retrieved document set consistently outperforms the sparse-retrieved documents from BM25. In Table 10, it is also observed that the intrinsic diversity from BM25-retrieved documents is also not as good as that from Contriever.
Dense retrievers like Contriever perform a semantic match between the query and document, whereas BM25 performs only a lexical match with no understanding of semantics. Picking semantically similar information is more important in SynthesizRR since we include a task inversion step, which can change the lexical structure of the text without changing its semantics significantly. Thus, we want contexts which are semantically related to Gold data, to which we can apply stylistic or formatting transformations using a task-inversion prompt to bring it closer to Gold. We might conclude that dense retrieval models are more suitable for SynthesizRR.
Appendix C Influence of retrieval corpus on domain shift
AG News($4$K)
Corpus DeBERTa ($\uparrow$) Mauve ($\uparrow$) Self-BLEU-5 ($\downarrow$) Entity Ent. ($\uparrow$)
RN/Dom 85.39 ± 0.8 92.58 0.23 6.72
RN/Rnd 35.57 ± 6.1 83.39 0.22 7.07
RN/Reg 84.17 ± 0.7 88.88 0.26 6.72
Hyperpartisan($2$K)
Corpus DeBERTa ($\uparrow$) Mauve ($\uparrow$) Self-BLEU-5 ($\downarrow$) Entity Ent. ($\uparrow$)
RN/Dom 78.77 ± 2.8 66.94 0.35 6.11
RN/Rnd 78.77 ± 3.5 61.45 0.25 7.40
RN/Reg 72.00 ± 2.0 65.59 0.35 6.12
Table 12: Effect of corpus-swapping for SynthesizRR 32-shot No-Retr-ICL. We generate only $4$k rows for AG News to reduce costs.
Our expectation is that SynthesizRR can flexibly specialize students to different domains by transparently changing the retrieval corpus, while keeping a frozen LLM. To quantify how changing the retrieval corpus might affect earlier metrics, we switch the news corpus for Hyperpartisan and AG News. We had assumed RealNews/Dom was the most suitable corpus (in-domain), and the others will cause domain-shift. In the following RQs, we validate the degree to which this assumption holds and the importance of information retrieval as the content sourcing mechanism in SynthesizRR.
RQ: Does modifying the corpus cause domain shift? Table 12 finds that the retrieval corpus highly influences the test performance (both student and intrinsic metrics). When grounding to a corpus with highly dissimilar entities (such as RealNews/Reg), all metrics drop significantly. Thus, we can conclude that an alternative content-source does indeed induce domain-shift. Mauve and distillation accuracy are highest for the in-domain corpus, while Self-BLEU and Entity entropy are highest for the random-retrieval results.
Figure 6: Retrieval counts for Hyperpartisan and AG News. The red dashed line represents the theoretical max, where all retrieved documents are unique. Note that the Random histogram plot is always 1 hence shows up as a straight line.
RQ: is retrieval essential for content sourcing? We measure the importance of retrieval by selecting top-k documents randomly from the in-domain corpus RealNews/Dom. We observe in Table 12 that retrieval using in-context learning queries plays a crucial role to the performance of AG News, as performance drops significantly in a random setting. Hyperpartisandoes not face such a drop. This matches our intuition in Table 1 that task-inversion is the more challenging step for Hyperpartisan, and a powerful LLM we can apply stylistic changes to most news articles. In both, Mauve suffers when entities no longer match Gold.
RQ: Do in-context queries retrieve redundant results? Figure 6 measures the overlap of top-50 retrieved documents from the 200 ICL queries, and finds that in most cases, the retrieved documents are unique, especially when using a large in-domain corpus. Thus, we can conclude that effective retrieval is important for the diversity of the synthetic dataset.
RQ: Can SynthesizRR work effectively with relatively small corpora? In our main results §section 5, we assumed the existence of a large corpus, with minimum size of 0.9M documents. As noted, this corpus need not be unlabelled examples for our task; we were able to successfully generate customer reviews and product questions for Humor, Categoryand Polarity tasks, while retrieving from a corpus of product information (title and description).
A potential problem with SynthesizRR is that corpuses of such massive size might be few in number. Thus, we compare the performance of SynthesizRR on CMU Movie Summary Bamman et al. (2013) and Yelp Dataset corpus Yelp (2022), which are between one to three orders of magnitude smaller than corpora used in Table 6. In Table 10 and Table 9, we see that SynthesizRR can perform suitably even with relatively small corpora (42k movie plots and 150k businesses). From the previous RQs, this suggests that the relevance of the corpus to the task is more important than the size of the corpus for the performance of SynthesizRR.
Appendix D Ablation: number of in-context examples
Figure 7: Left: DeBERTa-v3L test accuracy ()($\uparrow$)( ↑ ), center: entity entropy ()($\uparrow$)( ↑ ), right: Mauve ()($\uparrow$)( ↑ ) for SynthesizRR Retr-ICL. We vary the number of in-context examples from 0 to 8. Teacher LLMs LLaMa-2 Chat 13B and Claude Instant-v1 are compared on 6 tasks: AG News, Hyperpartisan, ToI Headlines, Category, Humor and Polarity. We do not report Category 8-shot due to model failures.
[…]
The use of in-context examples in the Retr-ICL variant of SynthesizRR leads to significant improvements in intrinsic and distillation metrics, as we saw in §5. Here, we do a deeper analysis on whether continually increasing the number of in-context examples yields a positive benefit.
In Figure 7 we look at the DeBERTa-v3L accuracy, entity entropy and MAUVE for our datasets with different numbers of in-context learning examples. We see that adding even a single in-context example can greatly increase the performance of all three metrics. However, no particular number of in-context examples consistently outperforms. For ClaudeV1, adding more in-context examples (up to 8) seems to always provide benefit, whereas with LLaMa2, we observe a peak and then reduction. Thus, the optimal number of in-context learning examples is a task dependent hyperparameter.
Figure 8 shows the lexical diversity i.e. Self-BLEU across datasets and number of in-context examples. As in §5 we observed that using in-context examples is neither positively nor negatively correlated with a lower Self-BLEU, despite using nucleus sampling with $p=0.9$. This may be because for all number of shots, task inversion is performed from a single source context and thus the generation does not divert significantly from the unique n-grams of the context. Thus we conclude that to affect lexical diversity, the number of in-context learning examples has no effect and we must instead focus on changing the retrieved contexts, perhaps by using a different retrieval model.
Appendix E Task inversion prompts and label verbalizations
Here we discuss the prompt templates and verbalizations that we use for the task inversion step for both FewGen and SynthesizRR. We use descriptive verbalizations as compared to the target label.
Additionally in the prompt, we place the retrieved document near the end, as prior work indicates that intermediate placements degrade LLM recall (Liu et al., 2023).
LLMs have a fixed window-size for conditional generation, so excessively long documents are truncated (from the end) up to $r_{max}=500$ tokens. This reserves the remaining window for in-context learning.
E.1 Hyperpartisan
Hyperpartisan is the task of detecting political bias in a news article. In transforming the retrieved news article article_retr[k] to one with such bias, typically there is the addition of mocking commentary and harsh political language which deeply criticizes the subject such as a person, policy or political event. On the other hand, articles in the opposite class gives a well-rounded opinion with a neutral tone. We include a length-attribute to ensure a long generation of one or two paragraphs.
Label Verbalization
true harsh political language, using a mocking tone and toxic commentary
false neutral language, using a reasonable tone and politically correct commentary
Table 13: Task-inversion verbalizations for Hyperpartisan.
Prompt E.1: Hyperpartisan FewGen In-context example:
“ Write a single news article using {label} . The written article should be 2 to 3 paragraphs long.
News Article: {icl[gold_text]}
Prompt:
“ Write a single news article using {label} . The written article should be 2 to 3 paragraphs long.
News Article: ”
Prompt E.2: Hyperpartisan SynthesizRR Retr-ICL In-context example:
“ News Article: {icl[article_retr]}
Rewrite the above news article using {label} . The rewritten article should be 2 to 3 paragraphs long.
Rewritten Article: {icl[gold_text]}
Prompt:
“ News Article: {article_retr[k]}
Rewrite the above news article using {label} . The rewritten article should be 2 to 3 paragraphs long.
Rewritten Article: ”
Prompt E.3: Hyperpartisan SynthesizRR No-Retr-ICL In-context example:
“ Rewritten Article: {icl[gold_text]}
Prompt:
“ News Article: {article_retr[k]}
Rewrite the above news article using {label} . The rewritten article should be 2 to 3 paragraphs long.
Rewritten Article: ”
E.2 ToI Headlines
ToI Headlines is a topic classification dataset of regional news headlines in India. Here we attempt to refine the retrieved news article by summarizing it into a short headline. We use verbalizations of the content of each class, as example generation here involves summarizing the content. We add an “India” location-attribute to guide the LLM generations to include regionalization to the Indian subcontinent. A length-attribute is included to restrict the length to one sentence.
Label Verbalization
sports sports in India
life-style health and lifestyle trends in India
education Indian examinations and education
entertainment the Indian entertainment industry
business business-related developments in India
city ongoing matters in any Indian city
environment environment-related events in Indian cities
tech technology news and the tech industry in India
elections elections and politics in India
world international news and events outside of India
Table 14: Task-inversion verbalizations for ToI Headlines.
Prompt E.4: ToI Headlines FewGen In-context example:
“ Write a headline for a news article about {label} . The headline should be a single sentence.
Headline: {icl[gold_text]}
Prompt:
“ Write a headline for a news article about {label} . The headline should be a single sentence.
Headline: ”
Prompt E.5: ToI Headlines SynthesizRR Retr-ICL In-context example:
“ News Article: {icl[article_retr]}
Write a headline for the above news article about {label} . The headline should be a single sentence.
Headline: {icl[gold_text]}
Prompt:
“ News Article: {article_retr[k]}
Write a headline for the above news article about {label} . The headline should be a single sentence.
Headline: ”
Prompt E.6: ToI Headlines SynthesizRR No-Retr-ICL In-context example:
“ Headline: {icl[article_retr]}
Prompt:
“ News Article: {article_retr[k]}
Write a headline for the above news article about {label} . The headline should be a single sentence.
Headline: ”
E.3 AG News
We consider task inversion for the AG News dataset to be generation of news summaries. We do not specify location modifiers as most Gold summaries are from US news. We add a length-attribute to restrict the output one or two sentences.
Label Verbalization
Business companies, industries, markets, trade, investments, entrepreneurship, economic policies, and other business-related developments
World international news, such as politics, diplomacy, conflicts, global events, international relations, human rights issues, and significant global trends
Sci/Tech scientific discoveries, technological advancements, innovations, research breakthroughs
Sports professional sports leagues, major tournaments, athletes, teams, match results, player transfers, coaching changes, sports-related controversies
Table 15: Task-inversion verbalizations for AG News.
Prompt E.7: AG News FewGen In-context example:
“ Write a summary for a news article about {label} . The summary should be one or two short sentences.
Summary: {icl[gold_text]}
Prompt:
“ Write a summary for a news article about {label} . The summary should be one or two short sentences.
Summary: ”
Prompt E.8: AG News SynthesizRR Retr-ICL In-context example:
“News Article: {icl[article_retr]}
Write a summary for the above news article about {label} . The summary should be one or two short sentences.
Summary: {icl[gold_text]}
Prompt:
“ News Article: {article_retr[k]}
Write a summary for the above news article about {label} . The summary should be one or two short sentences.
Summary: ”
Prompt E.9: AG News SynthesizRR No-Retr-ICL In-context example:
“ Summary: {icl[gold_text]}
Prompt:
“ News Article: {article_retr[k]}
Write a summary for the above news article about {label} . The summary should be one or two short sentences.
Summary: ”
E.4 Category
In the Category dataset, we determine the product category from a review written by a user. For task inversion in SynthesizRR we must retrieve a product and prompt the frozen LLM to generate a user review within the same product-category as the retrieval query. Thus, we include a style-attribute to allow minor typos in the generation and restrict to a few sentences using a length-attribute.
Label Verbalization
magazines magazines or periodicals covering various topics
camera_photo photography gear including cameras, lenses, accessories, or photo editing tools
office_products office supplies or equipment for professional and home office setups
kitchen kitchenware, appliances, or culinary tools for cooking and dining
cell_phones_service cell phone service accessories or service plans for communication and connectivity
computer_video_games computers, gaming consoles, video games, or related accessories
grocery_and_gourmet_food groceries, fruits and vegetables, gourmet treats, or specialty food items
tools_hardware tools, hardware, or equipment for DIY projects and home repairs
automotive auto parts, accessories, or tools for vehicle maintenance and enhancements
music_album music albums spanning various genres and artists
health_and_personal_care healthcare products, personal care items, or wellness essentials
electronics electronic devices, gadgets, personal tech, or home electronics
outdoor_living products for outdoor activities, gardening, or patio living
video movies, TV shows, and documentaries spanning various genres and artists
apparel clothing including casual wear, formal attire, seasonal outfits, activewear, or fashion accessories for men, women, and children
toys_games fun or educational toys and games for kids of all ages
sports_outdoors products for various sports and outdoor activities
books books in various genres and formats
software computer software for productivity or gaming covering either personal or professional needs
baby baby essentials, gear, or toys for infants and toddlers
musical_and_instruments musical instruments, accessories, or music production equipment
beauty beauty products, cosmetics, or skincare essentials, makeup, hair care, fragrances, or grooming essentials
jewelry_and_watches watches or jewelry pieces such as necklaces, bracelets, earrings, or rings, crafted in precious metals or adorned with gemstones for special occasions
Table 16: Task-inversion verbalizations for Category.
Prompt E.10: Category FewGen In-context example:
“ Write a product review about a product which is in the category of {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
“ Write a product review about a product which is in the category of {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
Prompt E.11: Category SynthesizRR Retr-ICL In-context example:
“ Product details: {icl[product_retr]}
Write a product review about the above product which is in the category of {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
“ Product details: {product_retr[k]}
Write a product review about the above product which is in the category of {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
Prompt E.12: Category SynthesizRR No-Retr-ICL In-context example:
“ Review: {icl[gold_text]}
Prompt:
“ Product details:
Write a product review about the above product which is in the category of {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
E.5 Humor
Asking humorous product questions is a challenge of the LLM’s task inversion capabilities, as it must generate a question which is funny from the retrieved product. Not all products have obvious humorous characteristics, thus the generation requires some ingenuity. We restrict the output to only the question to prevent explanations or extraneous product generations from the LLM.
Label Verbalization
humorous humorous
non_humorous solemn
Table 17: Task inversion verbalizations for Humor.
Prompt E.13: Humor FewGen In-context example:
“ Write a short {label} question about a product on Amazon. Only include the question.
Product Question: {icl[gold_text]}
Prompt:
“ Write a short {label} question about a product on Amazon. Only include the question.
Product Question: ”
Prompt E.14: Humor SynthesizRR Retr-ICL In-context example:
“ Product details: {icl[product_retr]}
Write a short {label} question about the above product on Amazon. Only include the question.
Product Question: {icl[gold_text]}
Prompt:
“ Product details: {product_retr[k]}
Write a short {label} question about the above product on Amazon. Only include the question.
Product Question: ”
Prompt E.15: Humor SynthesizRR No-Retr-ICL In-context example:
“ Product Question: {icl[gold_text]}
Prompt:
“ Product details: {product_retr[k]}
Write a short {label} question about the above product on Amazon. Only include the question.
Product Question: ”
E.6 Polarity
Polarity is a review-sentiment classification task. In SynthesizRR, the difficulty is increased as we must generate a review from a product. For task inversion, we prompt the LLM to generate a review which can have either positive or negative sentiment and include details from the retrieved product. As with Category, we allow typos and restrict the length to a few sentences using a length-attribute in the prompt.
Label Verbalization
positive what the reviewer liked about the product, how the reviewer found it easy to use the product, or the reviewer’s positive experience with the product
negative what the reviewer disliked about the product, how the reviewer found it challenging to use the product, or the reviewer’s negative experience with the product
Table 18: Task inversion verbalizations for Polarity.
Prompt E.16: Polarity FewGen In-context example:
“ Write a review about a product on Amazon which discusses {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
“ Write a review about a product on Amazon which discusses {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
Prompt E.17: Polarity SynthesizRR Retr-ICL In-context example:
“ Product details: {icl[product_retr]}
Write a review about the above product on Amazon which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
“ Product details: {product_retr[k]}
Write a review about the above product on Amazon which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
Prompt E.18: Polarity SynthesizRR No-Retr-ICL In-context example:
“ Review: {icl[gold_text]}
Prompt:
“ Product details: {product_retr[k]}
Write a review about the above product on Amazon which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: ”
Appendix F Example generations
Here we showcase examples from the best-performing SynthesizRR approach (3-shot No-Retr-ICL using LLaMa-2 Chat 13B) for each of our 6 tasks. For brevity, we do not show the ICL examples, only the retrieved article and generated text.
Class Example
(Retrieved document) News Article: Gov. Andrew Cuomo Wednesdy afternoon staged a press conference to blast federal immigration officials, threatening to sue if agents don’t stop what he described as "reckless and unconstitutional enforcement" practices. Immigration and Customs Enforcement Deputy Director Thomas D. Homan responded hours later with a six-paragraph statement that accused Cuomo of "grandstanding." Cuomo cited a raid last week on a farm in Rome, N.Y. Immigration and Customs Enforcement agents entered John Collins’ property and detained a worker, Marcial DeLeon Aguilar, without presenting a warrant to Collins. That worker was later arrested and is still being detained. "The reckless and unconstitutional practices ICE is deploying in our communities violate everything we believe in New York and are an assault on our democracy," Cuomo said in a prepared statement. "I demand ICE immediately cease and desist this pattern of conduct, and if they fail to do so, I will pursue all available legal recourse and commit to doing everything in my power to protect the rights and safety of all New Yorkers." At about 7:30 p.m., Homan issued this response: "As a native New Yorker who began my law enforcement career as a police officer in the state, I was disappointed to learn about the Governor’s grandstanding today over the issue of immigration enforcement. The Governor’s comments were inaccurate and an insult to ICE’s sworn law enforcement officers who conduct their lawful mission professionally and with integrity. These brave men and woman leave the safety of their homes every day to protect this great nation and our communities.
hyperpartisan What a joke! Governor Andrew Cuomo had the audacity to call out ICE agents for doing their jobs, and now he’s threatening to sue them? Give me a break! The nerve of this guy, talking about "reckless and unconstitutional enforcement" practices when his own state is infested with illegal immigrants who are committing crimes left and right. Cuomo is just trying to grandstand and pander to his liberal base, but the fact is, he’s failing to protect the citizens of New York. ICE agents are just doing their jobs, and they deserve respect for their hard work and dedication. Cuomo should be ashamed of himself for attacking these brave men and women. And let’s not forget, this is the same guy who is protecting child rapists and allowing them to roam free in his state. What a disgrace! Cuomo needs to focus on the real problems in his state, like the rampant crime and corruption, instead of trying to score political points by attacking ICE agents. It’s pathetic.
(Retrieved document) News Article: This week on "Watters’ World", Jesse Watters caught up with several House Democrats who boycotted President Trump’s inauguration, including Rep. John Lewis, the Atlanta Democrat who called the Republican’s election "not legitimate". Approaching Houston Congresswoman Sheila Jackson Lee, Watters asked why she was absent from the festivities: "I have a statement on record, thank you so very much," Jackson Lee said. WATCH: Watters Tracks Down Rep. John Lewis to Ask About Trump Criticism ’Watters’ World’ EXCLUSIVE: Jesse Visits Scene of Chicago Hate Crime Attack Watters Confronts Man Who Harassed Ivanka Trump on JFK Flight As Watters introduced himself as a Fox News reporter to Rep. John Conyers of Michigan, the Detroiter kept walking toward an open elevator, "no, sir, not right now." "I didn’t boycott it, I just didn’t go to it," Conyers said. Rep. Jerrold Nadler of New York gave a similar answer when asked, "for the reasons I stated publicly; they’re on the record." Asked if he was being divisive, the Manhattan Democrat said that if anyone wanted to unite the country, they would not have voted for Trump. Lewis told Watters he was late and couldn’t speak with him as he walked with a quick gait down a hallway. An aide obstructed Watters’ cameraman as the Congressman continued toward his destination.
non-hyperpartisan In a recent segment on "Watters’ World", Jesse Watters spoke with several members of the House of Representatives about their absence from President Trump’s inauguration. While some members provided brief statements, others declined to comment or did not respond to Watters’ inquiries. Rep. John Lewis, an Atlanta Democrat who had publicly questioned the legitimacy of Trump’s election, was approached by Watters but did not stop to speak with him. Lewis had previously stated that he would not attend the inauguration due to his disagreement with Trump’s presidency. Other members, including Rep. John Conyers of Michigan and Rep. Jerrold Nadler of New York, also declined to comment or did not respond to Watters’ questions. When asked why he did not attend the inauguration, Conyers simply stated that he "didn’t go to it". Nadler, who had publicly stated that he would not attend the inauguration due to his disagreement with Trump’s policies, reiterated that he did not support the President’s election. Watters’ segment sparked a discussion on the perceived divisiveness of the President and the responses of those who did not attend his inauguration. While some argued that the boycott was a form of protest against Trump’s policies and actions, others felt that it was a divisive move that further polarized the country. Regardless of one’s political beliefs, it is important to approach discussions and debates with respect and civility, and to strive for understanding and cooperation in the pursuit of common goals.
Table 19: Generated examples for Hyperpartisan task using SynthesizRR.
Class Example
(Retrieved document) News Article: New Delhi: First reports of exit poll indicate history in West Bengal and regime change in Kerala while it’s too close to call in Tamil Nadu. An exit poll conducted by Star News television channel has projected Congress-led United Progressive Alliance (UPA)’s victory in West Bengal and Kerala. The survey projects Trinamool Congress - Congress alliance wiping out more than three decades old Left Front government in West Bengal. According to the exit polls telecast in the channel on Tuesday says the TMC and Congress will win 181 and 40 respectively while Left Front will be reduced to 62 from 227. In Kerala, the Congress-led United Democratic Front is expected to win 88 while CPM-led Left Democratic Front will win 49. In Tamil Nadu, AIADMK is projected as winning 110 and ruling DMK will have 124 seats in the 234 member assembly. In a close contest, Tarun Gogoi led Congress has a chance to create history in Assam by returning to power for a third consecutive term,post poll surveys predict. The Headlines Today poll predicts 44 of the 126 seats for Congress in Assam,with potential ally AUDF slated to win 14.
elections Exit Polls predict regime change in Kerala, West Bengal; Tight contest in Tamil Nadu
(Retrieved document) News Article: India’s teen sensation P V Sindhu stormed into the third round of the World Badminton Championship with a straight game victory over Olga Golovanova of Russia in women’s singles match here yesterday Copenhagen: India’s teen sensation PV Sindhu stormed into the third round of the World Badminton Championship with a straight game victory over Olga Golovanova of Russia in women’s singles match here yesterday. PV Sindhu The men’s doubles pair of Manu Attri and Sumeeth Reddy B stunned 15th seeded Japanese duo of Hirokatsu Hashimoto and Noriyasu Hirata 21-19 21-19 in 44 minutes to advance to the third round. Sindhu, seeded 11th, took 40 minutes to prevail over her Russian opponent 21-12 21-17 in the second round match at the Ballerup Super Arena here. She will next take on sixth seeded Yeon Ju Bae of Korea. Sindhu won a total of 42 points as compared to 29 by the Russian girl. The world No. 12 from Hyderabad looked a bit rusty to start with and was initially trailing in the opening game. She was playing the catching-up game till 10-10 before Sindhu managed to reel off four consecutive points and surge ahead to 14-10. There was no looking back after that, as Sindhu swiftly sealed the game in her favour with Golovanova earning just two more points. In the second game, the Russian got her act together and opened up a big lead, moving up to 11-6 at the break.
sports 15-year-old PV Sindhu creates history, enters World Badminton Championship 3rd round
Table 20: Generated examples for ToI Headlines task using SynthesizRR.
Class Example
(Retrieved document) News Article: SANTIAGO Nov 14 A magnitude 6.1 earthquake hit central-north Chile on Wednesday, the U.S. Geological Survey said, causing buildings to sway in capital Santiago but causing no damage. The USGS said the quake struck on the coast, 54 miles (88 km) north of Coquimbo, at 4:02 p.m. (1902 GMT). It was initially reported as a magnitude 6.0 quake. Chile’s state emergency office, Onemi, said there were no reports of injuries, disruptions to basic services or infrastructure damage as a result of the tremor. The navy said the quake did not meet the conditions needed to trigger a tsunami. World No.1 copper producer Codelco said its operations were normal, and no mines in the metal-producing country reported damages. Seismically active Chile was battered in February 2010 by an 8.8 magnitude quake and ensuing tsunami, which caused billions in damage and killed hundreds.
World A magnitude 6.1 earthquake hit central-north Chile on Wednesday, causing buildings to sway in capital Santiago but causing no damage, the USGS said.
(Retrieved document) News Article: MONTREAL - Canadian astronaut David Saint-Jacques is bracing for some tense moments during next Monday’s launch of the Soyuz rocket that will send him and two others to the International Space Station. On Oct. 11, a rocket failure forced a Soyuz capsule carrying two astronauts to abort and make an emergency landing. Russia suspended all manned space launches pending an investigation before giving the green light Nov. 1. Saint Jacques spoke to reporters today from the launch site in Kazakhstan where he’s in quarantine. He said the most dangerous part of the six-month mission is the 10-minute Soyuz launch and the six hours that follow before docking. He said half of the last two-and-half years of training was dedicated to his role as Soyuz co-pilot for the trip to the space station. Saint-Jacques, 48, said once at the space station he’ll be able to focus on work and life aboard the outpost. The Canadian astronaut’s first space voyage, with NASA astronaut Anne McClain and Oleg Kononenkoof of the Russian space agency Roscosmos, was originally scheduled to launch Dec. 20. It was moved ahead after Russian authorities concluded their investigation into the failed launch.
Sci/Tech commentary Saint-Jacques says the Soyuz launch is the most dangerous part of the six-month mission but he’s been training for it for half of the last two-and-a-half years.
Table 21: Generated examples for AG News task using SynthesizRR.
Class Example
(Retrieved document) Product details: MSD ASY17296 Wiring Harness Factory Terminated MSD Ignition 6ALN (6430) Wiring Harness ASY17296 New harness with factory crimped connections a perfect item to have as an extra or replacement.
camera_photo I have a MSD Ignition 6ALN (6430) and this wiring harness is a great addition. The factory terminated connections are perfect and make installation a breeze. The harness is well made and will definitely last. I have been using it for a few weeks now and it has performed flawlessly. Highly recommend it to anyone looking for a high quality wiring harness.
(Retrieved document) Product details: Conair Travel Smart 3-Dial Combination Lock (6 Pack) Safely Locks Your Luggage With Tsa Recognized Locks Comes With A Red Inspection Status Feature Travel Sentry Approved
apparel I purchased this lock for a business trip to Europe. I’ve used it several times already and so far, so good. The TSA inspectors have been able to open it without a hitch, which is great!
Table 22: Generated examples for Category task using SynthesizRR.
Class Example
(Retrieved document) Product details: Keeping Up With The Kardashians - Season 5 Region E! brings viewers back into the sensational and completely unpredictable world of one of televisions favourite families with the fourth season of the hit series Keeping Up with the Kardashians. This season, Khlo and her husband, Los Angeles Lakers star, Lamar Odom, begin their new lives as a married couple. Meanwhile, Kourtney is pregnant and back together with Scott, and Kim and her boyfriend, NFL player Reggie Bush, work on their relationship. Through all this change, the Kardashian family must find ways to evolve and keep their close-knit ties.
humorous Does anyone else notice that this is the only show where you can hear people say things like, "OMG I am totally exhausted, I need to take a nap.? And it’s not considered weird?
(Retrieved document) Product details: Coleman Camper’s Storage Tote Securely organize, store and carry camping gear 23" x 14" x 17" Lightweight 600D polyester construction Multiple pockets and compartments Convenient carry strap The Coleman Camper’s Storage Tote is a great way to store, organize and carry your gear – use it in the garage, in the trunk or RV, and at the campsite. Constructed of reliable, lightweight 600D polyester, the Camper’s Storage measures 23" long, 14" wide and 17" tall. It features multiple pockets and compartments for organizing all your gear, and a convenient carry strap so it’s easy to bring along.
non_humorous is this product big enough for all the camping gear for a family of four?
Table 23: Generated examples for Humor task using SynthesizRR.
Class Example
(Retrieved document) Product details: Birth Of Soul 1 No half measures here, but a real statement of intent, particularly with that Volume One tag. 28 cuts that helped define that thing called soul - you can argue the point, but you can’t argue the music. Sublime.
positive This album was the beginning of my love affair with soul music. These tracks are timeless and still sound amazing today.
(Retrieved document) Product details: Beware! the Lies of Satan It’s high time we pull the covers off the enemy and expose him for who he really is, the murderer, destroyer, their, deceiver and biggest liar of all time. In this book , you will begin to discover the truth about our heavenly Father and how you can stand in victory over the devil.
negative The book does not live up to it’s promise. There is no revelation of truth about our heavenly father, or any insight into Satan’s lies. It is simply a polemic diatribe against Satan, with no concrete solutions to any of life’s problems.
Table 24: Generated examples for Polarity task using SynthesizRR.
Appendix G Data Preprocessing
G.1 Datasets
Other datasets are not pre-processed.
G.2 Corpora
  • RealNews: we use the article text field and download the data from https://github.com/rowanz/grover/tree/master/realnews.
  • RealNews/Regional is a subset of RealNews (Zellers et al., 2019). It includes 2.7M articles from non-US and non-EU websites. We manually check RealNews websites and identified 141 regional-news websites with headquarters in 56 non-US and non-EU countries: India, Pakistan, Nigeria, Philippines, etc.
  • RealNews/India is further filtered to only include Indian news websites.
  • RealNews/Dominant is the remaining 30.1M articles from 1063 news websites headquartered in 20 countries (of which over $75\%$ are US-based).
  • Products: We pull the data from https://nijianmo.github.io/amazon/index.html#complete-data and concatenate title and description
Appendix H Teacher and Student hyperparameters
H.1 Teacher LLM hyperparams
For LLaMa-2 Chat 13B, we use the implementation from HuggingFace: https://huggingface.co/TheBloke/Llama-2-13B-fp16 and run it at half-precision.
For Claude Instant-v1, we use Claude Instant v1.2: https://www.anthropic.com/news/releasing-claude-instant-1-2
We use a batch size of 1 for all generations as we have long contexts and encountered failures with higher batch sizes. We use nucleus sampling with top-p=0.9.
H.2 Student LM hyperparams
We use the same hyperparameters for DeBERTa-v3L and DistilBERT as (Yu et al., 2023a):
  • DistilBERT: Learning rate of 5e-5, gradient_accumulation_steps of 1, batch_size 32. We use the Adam optimizer with weight_decay of 1e-4 and epsilon of 1e-6. We use max_sequence_length of 512.
  • DeBERTa-v3L: Learning rate of 2e-5, gradient_accumulation_steps of 8, batch_size 4. We use the Adam optimizer with weight_decay of 1e-4 and epsilon of 1e-6. We use max_sequence_length of 512.
We train all students for 6 epochs. Following (Yu et al., 2023a), we use warmup for 6% of the training steps.
H.3 Oracle model hyperparams
To train the DeBERTa-v3-Large oracle model for Label Preservation, we use a grid search over 9 combinations: 3 learning rates {2e-5, 5e-5, 1e-4} by 3 batch-sizes {1, 4, 16} (with same graident accumulation). We train on 80% of the Gold training data and use the remaining 20% as validation.
H.4 Retriever
We use Contriever from HuggingFace library: https://huggingface.co/facebook/contriever.
We pass a batch-size of 512 for embedding.
Appendix I Computational budget
We run all our models on AWS Elastic Cloud Compute [^2] https://aws.amazon.com/ec2/ using 20 p3dn.24xlarge machines.
I.1 Information Retrieval
The corpora was embedded by us and the trivial was done using the Faiss library. [^3] https://faiss.ai/index.html We orchestrate 80 copies of Contriever using the Ray distributed framework [^4] https://docs.ray.io/en/latest/index.html to embed the RealNews and Products corpus in $\sim$ 3 hours each.
I.2 Dataset synthesis
For LLaMa-2 Chat 13B, we create 48 copies of the model and orchestrate it using the Ray distributed framework. Generation is done in roughly 6 hours per dataset of 8k rows.
To use Claude Instant-v1, we invoke AWS Bedrock [^5] https://docs.aws.amazon.com/pdfs/bedrock/latest/APIReference/bedrock-api.pdf using the boto3 library [^6] https://boto3.amazonaws.com/v1/documentation/api/latest/index.html. Generation from Claude was done at an AWS-account level RPM of 1600 and takes roughly 4 hours per dataset on 8k rows.
I.3 Student distillation
Each DeBERTa-v3-Large student model trains for between 1 to 3 hours on a single V100 GPU on 8k rows. Each DistilBERT student model trains in 1 hour to generate the data-map.
Appendix J Licensing
We use datasets that have been released in prior work with various open licenses. Specifically:
J.1 Datasets
J.2 Corpora