### v3

# SynthesizRR: Generating Diverse Datasets with Retrieval Augmentation

Abstract
It is often desirable to distill the capabilities of large language models (LLMs) into smaller student models due to compute and memory constraints. One way to do this for classification tasks is via dataset synthesis, which can be accomplished by generating examples of each label from the LLM. Prior approaches to synthesis use few-shot prompting, which relies on the LLM’s parametric knowledge to generate usable examples. However, this leads to issues of repetition, bias towards popular entities, and stylistic differences from human text. In this work, we propose Synthesize by Retrieval and Refinement (SynthesizRR), which uses retrieval augmentation to introduce variety into the dataset synthesis process: as retrieved passages vary, the LLM is “seeded” with different content to generate its examples. We empirically study the synthesis of six datasets, covering topic classification, sentiment analysis, tone detection, and humor, requiring complex synthesis strategies. We find that SynthesizRR [^1] https://github.com/amazon-science/synthesizrr greatly improves lexical and semantic diversity, similarity to human-written text, and distillation performance, when compared to 32-shot prompting and four prior approaches.
SynthesizRR: Generating Diverse Datasets with Retrieval Augmentation
[…]
1 Introduction
Figure 1: Synthetic examples from few-shot generation (middle) and SynthesizRR (bottom). Our approach incorporates a content sourcing step which retrieves documents from a corpus: for the task of detecting political bias, a news article is retrieved and the teacher LLM is prompted to produce a biased version. The resulting synthesis procedure yields diverse examples which more closely match human-written examples.
Figure 2: Abstract depiction of the SynthesizRR procedure. In the content sourcing stage, we retrieve $K$ unique document $\{r_{1},\dots,r_{K}\}$ from a large corpus for each in-context covariate $x_{\text{ICL}}$. The task-inversion stage of synthesis uses a parameterized context refinement prompt $\mathcal{P}_{\tau}$, which takes parameters $R_{inv}$ (inversion instruction), $r_{k}$ (a retrieved document), and $\mathcal{V}(y_{\text{ICL}})$ (the verbalized target label). A generalist teacher LLM autoregressively generates a synthetic covariate. Each in-context example thus produces $K$ unique synthetic examples $\{\tilde{x}_{1},\dots,\tilde{x}_{K}\}$, which we include in the dataset with target $y_{\text{ICL}}$.
Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023; Bubeck et al., 2023), LLaMa (Touvron et al., 2023b) and Claude (Bai et al., 2022) are versatile generalist models, capable of solving multiple tasks without parameter tuning via zero-shot or few-shot prompting. In comparison, previous approaches fine-tuned variants of BERT (Devlin et al., 2019) on task-specific demonstrations, producing specialist models. These smaller specialist models are more economical at inference time, but require at least thousands of examples to train.
Recent work has sought to avoid this reliance on manually created examples by fine-tuning specialist models on synthetic datasets via teacher-student distillation (West et al., 2022). This has applications in classification (Yu et al., 2023a; Ye et al., 2022a, b), human-preference alignment (Lee et al., 2023; Bai et al., 2022), language understanding (Meng et al., 2022; Schick and Schütze, 2021), and even tabular data (Borisov et al., 2022). However, synthetic data has limitations. As Yu et al. (2023a) note, naive prompts generate texts with limited diversity and reflecting biases of the teacher LLMs.
Figure 1 illustrates the few-shot synthesis approach (Ye et al., 2022a, b; Yehudai et al., 2024a), which we refer to as FewGen, for the task of detecting politically-biased articles. With a suitable prompt and in-context examples, sampling continuations from an LLM generates plausible news in the biased style we seek to detect. However, as thousands of completions are sampled from a fixed prompt, we observe repetition, bias towards popular entities, and stylistic differences from human-written texts. Specialist models distilled from such low diversity datasets may not learn the task well.
In this work, we seek to alleviate the lack of diversity in synthetic data. We suggest that dataset synthesis may be decomposed as two distinct LLM competencies: content sourcing, where the LLM obtains relevant information for the task, and task inversion, where the LLM generates a synthetic input using a target-conditioned prompt. Prior work has focused mainly on task inversion, while implicitly using the LLM’s parametric memory for content sourcing. In contrast, we investigate the importance of an explicit content sourcing stage.
We propose Synthesize by Retrieval and Refinement (SynthesizRR), an example synthesis procedure guided by a retrieval corpus. In the content sourcing step, we use in-context learning covariates as retrieval queries to extract dozens of documents per query from a domain-specific corpus. Subsequently, a generalist LLM performs task inversion on each retrieved document. As each prompt uses a unique retrieved document, our synthesis procedure generates diverse examples, enriched with a broad spectrum of real-world entities and assertions.
We benchmark SynthesizRR against FewGen on six text classification tasks, selected carefully to measure a variety of different styles of dataset synthesis. Our experiments (§5) reveal that SynthesizRR significantly surpasses FewGen in diversity and resemblance to human-authored texts, even though both procedures utilize the same frozen LLM. In §6, we see that student classifiers fine-tuned on SynthesizRR-generated data perform better than those fine-tuned on FewGen. Finally, in §7, we compare SynthesizRR to four state of the art approaches for synthesis of classification datasets, and find SynthesizRR gives higher diversity datasets, better matching human-written instances, and leads to higher student accuracy in most cases.
Our contributions are as follows: (1) we propose a new method of example synthesis for teacher-student distillation, which grounds the task inversion step using a retrieval corpus; (2) we introduce the SynthesizRR RetrICL algorithm to create a realistic in-context learning set for our method; (3) we empirically analyze the synthesis of six challenging classification tasks, comparing our method’s textual diversity and similarity and downstream task accuracy to existing approaches; (4) we pinpoint factors affecting the quality of our synthetic datasets by varying the amount of supervised data, corpus relevance to task, number of in-context examples, and sparse vs. dense retrieval.
2 Background and Task setup
In this paper, we focus on generating datasets for challenging text classification tasks. Denote an example as consisting of input text $x$ and output $y\in\mathcal{Y}$ for output space $\mathcal{Y}$ of $C$ classes. Our goal is to produce a synthetic dataset $\mathbin{\mathbin{\mathcal{D}_{\textsc{Synth}}}=\left\{(\mathbin{\tilde{x}^{i}},\mathbin{y^{i}})\right\}^{m}_{i=1}}$ and train a specialist language model $\mathcal{M}_{\text{S}}$ (e.g. a BERT-style pre-trained model (Devlin et al., 2019)). We create $\mathbin{\mathcal{D}_{\textsc{Synth}}}$ via task inversion: repeatedly prompting a teacher language model $\mathcal{M}_{\text{LM}}$ to generate synthetic covariates $\tilde{x}$ given corresponding labels $y$. We denote the student’s task (predicting $y$ from $x$) as $\tau$ and the teacher’s task (generating $x$ given $y$) as $\tau_{inv}$.
SynthesizRR aims to address the lack of diversity by leveraging retrieval during the content sourcing step. We assume the existence of a corpus $\mathcal{R}$ where each document may hold task-relevant information. However, documents need not originate from the same distribution as our task covariates; even distantly related documents can yield valuable synthetic examples. For instance, we shows that we can successfully generate reviews and humorous questions from a corpus of product descriptions. We also assume access to a seed set of examples $\mathcal{D}_{\textsc{Seed}}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ which is sufficiently large to represent the classes but small enough to be manually compiled by a user in a few hours; in experiments, we use the in-context learning set as $\mathcal{D}_{\textsc{Seed}}$. Importantly, we assume the seed set is insufficient to train an effective student, and a larger $\mathbin{\mathcal{D}_{\textsc{Synth}}}$ ($m>>n$) is needed.
Figure 2 illustrates our method for generating distributionally similar covariates. Initially, we retrieve documents based on the examples in $\mathcal{D}_{\textsc{Seed}}$, assuming that the corpus contains sufficient domain-similar documents. We then construct a context refinement instruction to perform task inversion on each retrieved document. This approach provides the LLM with a unique and grounded prompt for each generated example, thereby circumventing the need for the teacher LLM to memorize extensive corpus data within its limited parameters. Task inversion may be challenging due to the mismatch between retrieved documents and test examples; to overcome this, we limit our investigation to teacher LLMs demonstrating strong instruction-following capabilities (Ouyang et al., 2022; Touvron et al., 2023b; Bai et al., 2022).
Algorithm 1 SynthesizRR RetrICL
Input A set of seed examples $\mathcal{D}_{\textsc{Seed}}$, retrieval corpus =$\mathcal{R}=\{r_{k}\}$, retrieval model $\mathcal{M}_{\mathrm{ret}}$, expansion factor $K$, cosine-similarity criterion $(s_{\alpha},s_{\beta})$, teacher model $\mathcal{M}_{\text{LM}}$, prompt template $\mathcal{P}_{\tau}$, context refinement instruction $R_{inv}$, verbalizer $\mathcal{V}:\{y_{1},\dots,y_{C}\}\rightarrow\{v_{1},\dots,v_{C}\}$.
Output Synthetic dataset $\mathbin{\mathcal{D}_{\textsc{Synth}}}{}$
Procedure $\textsc{SynthesizRR}(\mathcal{D}_{\textsc{Seed}},\mathcal{R})$:
$\mathcal{D}_{\textsc{Retr}}\leftarrow\emptyset$
$\mathcal{D}_{\textsc{ICL}}\leftarrow\emptyset$
$\mathbin{\mathcal{D}_{\textsc{Synth}}}{}\leftarrow\emptyset$
$\triangleright$ Content sourcing using retrieval:
for $(x,y)\in\mathcal{D}_{\textsc{Seed}}$ do
     $[r_{1},\ldots,r_{K}]\leftarrow\mathcal{M}_{\mathrm{ret}}(x)$
     $\Gamma_{K}\leftarrow[r_{1},\ldots,r_{K}]$
     $\mathcal{D}_{\textsc{Retr}}\leftarrow\mathcal{D}_{\textsc{Retr}}\cup\{(x,y,\Gamma_{K})\}$
$\triangleright$ In-context learning set construction:
for $(x,y,\Gamma_{K})\in\mathcal{D}_{\textsc{Retr}}$ do
     for $r_{k}\in\Gamma_{K}$ do
         $\mathcal{D}_{\textsc{ICL}}\leftarrow\mathcal{D}_{\textsc{ICL}}\cup\{(r_{k},x)\}$ if $s_{\alpha}\leq\cos(x,r_{k})\leq s_{\beta}$
$\triangleright$ Task inversion:
for $(x,y,\Gamma_{K})\in\mathcal{D}_{\textsc{Retr}}$ do
     for $r_{k}\in\Gamma_{K}$ do
[…]
              $\mathbin{\tilde{x}}^{i}_{j}{}{\sim}{}\mathcal{M}_{\text{LM}}\left(\cdot|\mathbin{\tilde{x}}^{i}_{<j},\mathcal{P}_{\tau}(R_{inv},r_{k},\mathcal{V}(y)),\mathcal{D}_{\textsc{shots}}\right)$
         $\mathbin{\mathcal{D}_{\textsc{Synth}}}{}\leftarrow\mathbin{\mathcal{D}_{\textsc{Synth}}}{}\cup\{(\tilde{x}^{i},y)\}$
return $\mathbin{\mathcal{D}_{\textsc{Synth}}}{}$
3 Method
Algorithm 1 shows our dataset generation method. We distill a student model in these steps:
Step 1. Content sourcing using retrieval: SynthesizRR uses each in-context covariate $x_{{}_{\text{ICL}}}$ as a query for information retrieval, in addition to its subsequently role during in-context learning. For each query, we retrieve $K$ documents $\Gamma_{K}=[r_{1},\dots,r_{K}]$ of progressively decreasing cosing similarity using the dense retriever $\mathcal{M}\mathrm{ret}$. We retain documents with cosine similarity in (0.4, 0.9), to ensure minimum similarity while excluding overly similar documents as potential duplicates of $x_{{}_{\text{ICL}}}$. Each resulting triplet $(x_{{}_{\text{ICL}}},y_{{}_{\text{ICL}}},\Gamma_{K})$ is appended to set $\mathcal{D}_{\textsc{Retr}}$.
Step 2. In-context set construction: The subsequent task inversion step also benefits from in-context demonstrations, but it is challenging to construct demonstrations which effectively captures our context refinement task $r_{k}^{i}\rightarrow\mathbin{\tilde{x}}^{i}$. We explored two approaches to in-context learning.
1. RetrICL: we use retrieval to construct a set of ICL examples $\mathcal{D}_{\textsc{ICL}}$, such that each ICL example mirrors the format of our task-inversion prompts. We select top-1 and top-2 retrieved results from the densely retrieved results, and use a cosine-similarity criterion $s_{\alpha}\leq\cos(x_{{}_{\text{ICL}}},r_{k})\leq s_{\beta}$ to asses the potential match between the retrieved document $r_{k}$ and $x_{{}_{\text{ICL}}}$. Although the in-context pair may not match exactly, they demonstrate the required format as per Appendix G.
2. Non-RetrICL: a baseline method, which uses retrieval for content sourcing, but not for in-context learning. For each generation we select $N=32$ ICL examples at random from $\mathcal{D}_{\textsc{Seed}}$. Each example is appended with a prefix like “News Article:” or “Product details:” but we do not add the context refinement instruction. After the ICL examples, we append the retrieved document $r_{k}$ and context refinement instruction $R_{inv}$ to form the final prompt. This format closely mirrors the in-context learning prompt used by FewGen, but also incorporates content-sourcing elements $r_{k}$ and $R_{inv}$. This baseline highlights the value added by constructing $\mathcal{D}_{\textsc{ICL}}$ in the RetrICL approach.
Step 3. Task inversion using context refinement: The minimum elements of a task inversion prompt $\mathcal{P_{\tau}}$ are the context refinement instruction $\mathcal{I}_{inv}$ and target $y$. We use a verbalizer function $\mathcal{V}$ (Schick and Schütze, 2021; van de Kar et al., 2022) to provide a unique text representation of each label, i.e. $\mathcal{V}:\mathcal{Y}\rightarrow\{v_{1},\dots,v_{C}\}$. We follow prior work on classification-based task inversion (Schick and Schütze, 2021; Ye et al., 2022a, b; Yu et al., 2023b; Gao et al., 2023) and use descriptive verbalizations to induce label-separability in the final dataset.
FewGen uses the standard causal language modeling objective to induce next-token probabilities from teacher LLM, $\mathcal{M}_{\text{LM}}$. Nucleus sampling Holtzman et al. (2019) is used to autoregressively sample next tokens until the <eos> token is generated. This becomes synthetic example $\mathbin{\tilde{x}}^{i}$.
[…]
For each label $y$, we fix this prompt and sample $m/C$ times to generate the synthetic dataset.
In SynthesizRR, we create the synthetic dataset from each triplet in $\mathcal{D}_{\textsc{Retr}}$. The retrieved documents $\Gamma_{K}=[r_{1},\dots,r_{K}]$ have lexical and semantic overlap with the query $x_{{}_{\text{ICL}}}$. However, corpus documents may be distributionally dissimilar from real task covariates, due to the nature of documents or chunking process (Mialon et al., 2023). To address this, we use $\mathcal{M}_{\text{LM}}$ to perform task inversion from the content of each retrieved document, a process we refer to as contextual refinement. $\mathcal{P_{\tau}}$ is thus composed from the contextual refinement instruction $\mathcal{R}_{inv}$, each document $r_{k}\in\Gamma_{K}$, and the verbalized target for the query, i.e. $\mathcal{V}(y_{{}_{ICL}})$. The LLM’s context window thus sees a unique and grounded prompt when auto-regressively generating each synthetic input $\mathbin{\tilde{x}}^{i}$:
[…]
for all documents $r_{k}\in\Gamma_{K}$. We continue to use nucleus sampling to get diverse generations. Each original in-context example thus produces $K$ unique synthetic examples $\{\tilde{x}_{1},\dots,\tilde{x}_{K}\}$; we call $K$ the “expansion factor”. To promote adherence to $\mathcal{R}_{inv}$, we sample pairs from $\mathcal{D}_{\textsc{ICL}}$ to create in-context examples following the same format. Our final dataset is constructed as:
$\mathcal{D}_{\textsc{Synth}}=\bigcup\limits_{(x,y,\Gamma_{K})\in\mathcal{D}_{\textsc{Retr}}}{}\bigcup\limits_{r_{k}\in\Gamma_{K}}\left\{(\mathbin{\tilde{x}^{i}},y)\right\}$.
Step 4. Student distillation: The student is fine-tuned on $\mathbin{\mathcal{D}_{\textsc{Synth}}}$ by passing the BERT [CLS] token embedding of $\mathbin{\tilde{x}}$ through a feedforward layer. This produces a probability distribution over the label space $C$. We optimize the cross-entropy loss of the true label $y$. As we derive $\mathbin{\tilde{x}}$ from a teacher LLM, this can be considered a form of symbolic knowledge distillation (West et al., 2022).
Dataset Class Train, Test Corpus Difficulty
AG News 4 $115\text{k},7.6\text{k}$ RN/Dom Easy
ToI Headlines 10 $52\text{k},10\text{k}$ RN/Ind Easy
Hyperpartisan 2 $516,65$ RN/Dom Medium
Polarity 2 $72\text{k}^{*},7.2\text{k}^{*}$ Products Medium
Category 23 $30\text{k}^{*},2.4\text{k}^{*}$ Products Medium
Humor 2 $15\text{k},3\text{k}$ Products Hard
IMDb 2 $20\text{k},25\text{k}$ Movies Medium
SST-2 2 $54\text{k},872$ Movies Medium
Table 1: Dataset statistics and our estimate of task inversion difficulty. Downsampled for convenience.
4 Experimental Setup
Tasks and their difficulty. We perform our main experiments on the first 6 datasets in Table 1, selected carefully to measure how the teacher LLM performs on task inversion tasks of varying difficulty. Previous work only benchmarked sentiment and topic classification datasets like IMDb Maas et al. (2011) and AG News (Zhang et al., 2015). We broaden from topic classification, which primarily involves summarization during the task inversion step, which LLMs are adept at (Goyal et al., 2022). Hyperpartisan (Kiesel et al., 2019) detects bias in political news, so the task inversion step includes a more substantial rewriting of neutral retrieved articles to form biased examples. Category and Polarity are prevalent product review tasks (Yu et al., 2023a, b; Gao et al., 2023); we generate reviews from retrieved products which must conform to categorical and sentiment classes. Task inversion for Humor (Ziser et al., 2020) involves generating humorous questions from retrieved product details, which requires additional skills from the teacher. Prompts for all tasks are in Appendix G.
Corpus Domain Size Doc. Tokens
[…]
Table 2: Corpus statistics with LLaMa2 tokenizer.
Table 2 describes corpora used for retrieval. We consider five corpora in different domains, each with varying numbers of records. Three are subsets of RealNews (Zellers et al., 2019), as described in Appendix I: RealNews/Dominant (US/EU News), RealNews/Regional (Regional News), RealNews/India (Indian News). We also use Products (Amazon products metadata, (Ni et al., 2019)) and Movie Summary (movie summaries, (Bamman et al., 2013). Each task in Table 1 is associated with the corpus we consider most relevant. In §7, we compare to four prior approaches on three other tasks: IMDb Maas et al. (2011), SST-2 Socher et al. (2013) and AG News. These sentiment and topic tasks are less aligned with our goals and thus excluded from our main evaluation.
Method Example
Gold There is decent bass, but the highs are a bit soft. A quick tweak to my equalizer, and they’re great. After reading several of the reviews on Amazon, I was a bit worried about the sound, but now that I have them I’m very happy. They’re a good price, and sooooo much better than the little ipod-like earbuds I’ve tried before. Those never stayed in my ear, and the bass never made me happy.
FewGen I’ve been a very happy customer of this company for a long time. It is fast and does everything I need it to. I would definitely recommend it to anyone looking for a good external drive. However, I do have one issue with the product. The instructions that come with it are not very clear and I had a hard time figuring out how to properly use it.
(Retrieved Product) Portable Laptop Microphone. Connects to 1/8" mini microphone input on laptop. Right-angle shaped. Flat-frequency response.
SynthesizRR The portable laptop microphone is right-angled and has a flat-frequency response, making it easy to use for online meetings and interviews. It connects to the 1/8" mini microphone input on my laptop and has worked great for the past two months, but I have noticed some distortion in the audio when I move around too much. Overall, it’s a great value for the price and has made my remote work and video conferencing much more productive and efficient.
Table 3: Real and synthetic examples from “electronics” class of Category. Grey text indicates lack of specifics.
Models. We use Contriever (Izacard et al., 2022) for dense retrieval from each corpus. This performs a semantic match between the query and each document using cosine-similarity. In Appendix E, we also perform an ablation study using BM25 as a sparse retriever, which does lexical matching between each query-document pair.
As teacher models, we primarily use a frozen Llama-2 Chat 13B (Touvron et al., 2023b) for the task inversion step in SynthesizRR and FewGen. We also experiment with Claude Instant-v1 as described in Appendix J. For in-context learning (ICL) Brown et al. (2020), we select examples randomly from the train set: 50 ICL examples/class for multi-class and 100/class for binary tasks. We believe this is a realistic number of examples that a system designer could source if they were to put some effort into building a specialist model. We explore approaches to bootstrap this seed set in limited-supervision settings Appendix C.
Specialization performance is measured on student LMs DeBERTa-v3-Large (435M params, He et al. (2021)) and DistilBERT (66M params, Sanh et al. (2019)).
Figure 3: Self-BLEU ()($\downarrow$)( ↓ ) for ngrams n=1-5. Comparison: Gold, FewGen 0-shot, FewGen 32-shot, SynthesizRR 0-shot, SynthesizRR 3-shot RetrICL, SynthesizRR 32-shot Non-RetrICL.
Evaluation criteria. Text generation can be challenging to evaluate objectively in multi-task scenarios (Chang et al., 2024). Therefore in §5 we evaluate synthetic text based on several criterion, to detect behaviours we observe during synthesis as in Table 3. Self-BLEU (Papineni et al., 2002; Zhu et al., 2018) measures lexical diversity of the dataset based on $n$-gram overlap between pairs of examples. Entity entropy measures the diversity of entities using the probability distribution of each of 16 entity-types, inferred using spaCy’s en_core_web_lg (Honnibal et al., 2020). Datasets which over-represent popular entities score lower on entropy. On the other hand, Entity recall and Entity KL divergence compares the similarity of entities compared to Gold, and datasets which reproduce entities frequently seen in Gold data score higher. MAUVE (Liu et al., 2021) measures similarity to human-written text by using pretrained representations from a gpt2-xl model, indicating distributional differences in the generated text.
Figure 4: Entity entropy ()($\uparrow$)( ↑ ) on ToI (headlines) and Category (reviews). Comparison: Gold, FewGen 32-shot, SynthesizRR 3-shot RetrICL and SynthesizRR 32-shot Non-RetrICL. Zero-shot results are similar for SynthesizRR and worse for FewGen; we omit them.
Method Norp Org Person Gpe Recall ()($\uparrow$)( ↑ ) KL div. ()($\downarrow$)( ↓ )
Unique Entities
Gold 319 3943 3952 712 - -
FewGen* 43 480 400 73 0.05 -
SynzthRR 137 2718 1528 238 0.12 -
SynzthRR 109 1755 1012 178 0.10 -
Total Entities
Gold 843 7233 6096 1558 - -
FewGen* 94 775 506 96 0.23 3.10
SynzthRR 319 3991 1989 397 0.35 2.35
SynzthRR 314 2699 1464 363 0.32 2.52
Table 4: Entity similarity in Category ($8$K). We show the counts of unique and total entities for 4 entity-types. Entity recall measures the fraction of Gold entities co-occuring in the synthetic data; in the bottom half, we additionally weigh each entity by its frequency in Gold. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
5 Results: Intrinsic Evaluation
In this section, we focus on evaluating intrinsic properties of the generated datasets, including their diversity and entity coverage. We focus on a LLaMa-2 Chat 13B teacher LLM, retrieving from Contriever using corpora per Table 1 (we analyze changing the retrieval corpus in  Appendix D). We generate datasets of size in relation to the number of Gold rows: $8$K rows (AG News, ToI Headlines, Category), $4$K rows (Polarity) or $2$K rows (Hyperpartisan, Humor). Example generations are in Appendix H.
RQ: Does retrieval augmentation improve lexical diversity? Figure 3 shows lexical diversity within the dataset. Human-written texts (Gold) score high on lexical diversity (low Self-BLEU). FewGen texts tend to reuse the same words and phrases, leading to repeated text across generations (high Self-BLEU). SynthesizRR text has lexical diversity approaching human text for all n-gram values. We note in-context learning has an inconsistent effect; it improves the lexical diversity for news corpora but not for products.
RQ: Does SynthesizRR address entity diversity? Popularity bias is a phenomenon wherein LLM generations tend to over-represent popular “head” entities. This has been studied for QA tasks (Mallen et al., 2023; Kandpal et al., 2023).
In Figure 4 we see how SynthesizRR eliminates popularity bias across entity types. By sourcing from the long-tail of retrieval results ($k=50$), the generated dataset has much higher entity entropy compared to FewGen. This positions SynthesizRR closer to Gold, which also shows high entity entropy.
Method AG. Hyp. ToI Cat. Hum. Pol.
(Dataset size) ($8$K) ($2$K) ($8$K) ($8$K) ($2$K) ($4$K)
Zero shot
FewGen $56.57$ $53.68$ $62.79$ 63.2 $75.6$ $62.82$
SynzthRR 90.3 59.2 63.0 $61.06$ 82.9 78.6
Few shot
FewGen* $56.7$ $65.39$ $60.33$ $65.8$ $78.14$ $69.21$
SynzthRR 92.0 72.8 87.9 75.2 87.5 89.9
SynzthRR $91.76$ $67.86$ $67.16$ $75.14$ $86.95$ $83.2$
Table 5: MAUVE similarity score ($\uparrow$) using GPT2-XL embeddings. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
Method Teacher LM AG. Hyper. ToI Categ. Humor Polar. Avg
(Dataset size) ($8$K) ($2$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold - 91.015 78991.01578991.015\,78991.015 789 93.230 76993.23076993.230\,76993.230 769 $82.488$ 81.516 66781.51666781.516\,66781.516 667 $93.06$ $95.2645$ 89.43
Seed - $83.9$ $82.5$ $67.5$ $71.7$ $85.0$ $90.9$ 80.25
Zero-shot
FewGen LLaMa2 69.457 89569.45789569.457\,89569.457 895 72.6 $32.09$ $62.35$ 74.433 33374.43333374.433\,33374.433 333 $80.996$ 65.32
FewGen ClaudeV1 75.031 57975.03157975.031\,57975.031 579 57.538 46257.53846257.538\,46257.538 462 $23.298$ $47.1$ 49.866 66749.86666749.866\,66749.866 667 $87.481$ 56.72
SynthesizRR LLaMa2 83.510 52683.51052683.510\,52683.510 526 69.846 15469.84615469.846\,15469.846 154 74.4 68.9 82.5 $84.707$ 77.32
SynthesizRR ClaudeV1 83.9 72.307 69272.30769272.307\,69272.307 692 $71.83$ 66.783 33366.78333366.783\,33366.783 333 62.113 33362.11333362.113\,33362.113 333 88.7 74.29
Few-shot
FewGen* LLaMa2 $84.2$ 74.461 53874.46153874.461\,53874.461 538 73.7 68.641 66768.64166768.641\,66768.641 667 $88.38$ $90.899$ 80.05
FewGen* ClaudeV1 75.852 63275.85263275.852\,63275.852 632 58.461 53858.46153858.461\,53858.461 538 $72.224$ 68.833 33368.83333368.833\,33368.833 333 $82.94$ $91.2435$ 74.93
SynthesizRR LLaMa2 82.992 10582.99210582.992\,10582.992 105 78.461 53878.46153878.461\,53878.461 538 $73.262$ 72.4 90.2 $91.0015$ 81.38
SynthesizRR LLaMa2 85.2 79.1 $72.826$ 71.941 66771.94166771.941\,66771.941 667 88.773 33388.77333388.773\,33388.773 333 $88.197$ 81.00
SynthesizRR ClaudeV1 $83.7394$ 72.307 69272.30769272.307\,69272.307 692 $72.832$ 65.408 33365.40833365.408\,33365.408 333 83.393 33383.39333383.393\,33383.393 333 91.3 78.16
SynthesizRR ClaudeV1 83.710 52683.71052683.710\,52683.710 526 $72.00$ $72.45$ 67.816 66767.81666767.816\,66767.816 667 $76.2066$ $87.87$ 76.68
Table 6: Test Accuracy ($\uparrow$) after distilling DeBERTa-v3-Large student from LLaMa-2 Chat 13B and Claude Instant-v1. Contriever was used as the retriever in SynthesizRR. We report the average of 5 runs and rerun in cases where  std. dev. $\geq$6(indicating one or more models failed to converge). The top half considers zero-shot synthesis and bottom half uses in-context learning, and we bold the best result under each paradigm. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
RQ: How is entity similarity in synthetic data affected by grounding to an in-domain corpus? For the Category task we generate $8$K product reviews and randomly select $8$K Gold examples. In Table 4, we measure entity recall, and find that the occurrence of Gold entities is 100%-140% higher in SynthesizRR than FewGen. The KL divergence of each entity distribution is also lower. We finally consider the entity coverage (unique entities) and entity density (total entities). Compared to Gold, FewGen tends to produce fewer unique entities (places, events, languages, currencies, etc). Each FewGen example also has a lower density of entities, as visible in Table 3. SynthesizRR coverage and density more closely match Gold.
RQ: How distributionally similar are our generated examples and human-written examples? We see from MAUVE scores in Table 5 that zero-shot generations are quite dissimilar in both approaches compared to few-shot methods. Surprisingly, SynthesizRR generations are much more similar to human text than FewGen, despite the fact that nothing in our content sourcing strategy explicitly guides SynthesizRR generations to match the distribution of Gold.
We thus manually inspect generations and discover an interesting pattern which can be attributed to content sourcing. As shown earlier, and in Table 3, the density of entities is higher under SynthesizRR. FewGen produces generations which obey the prompt, but are very bland and do not include specifics. On the other hand, by obtaining information-rich documents, SynthesizRR is able to ground the task inversion step in details of the retrieved article/product. We hypothesise that this improves the MAUVE score towards Gold, which is similarly grounded in specifics.
6 Results: Student distillation
We have established that SynthesizRR generates more diverse datasets compared to a baseline approach. Now, we return to the application of training a specialist model based on these datasets.
Table 6 shows the results of training a DeBERTa-v3-Large student on datasets generated by SynthesizRR and FewGen, as well as baselines of tuning on the Gold set and Seed set. In the zero-shot setting, we find that SynthesizRR performs much better than FewGen, despite using the same frozen teacher LLM. Note that SynthesizRR uses in-context examples for retrieval here whereas FewGen does not; our method has some additional supervision here. However, in this setting, we see clear gains during the task inversion stage ($\uparrow$12for LLaMa and $\uparrow$17.6for Claude). Thus, having access to retrieval yields a better final dataset, almost on par with 32-shot FewGen.
With ICL, 3-shot SynthesizRR using the RetrICL strategy trains better students than 32-shot FewGen ($\uparrow$1.3for LLaMa and $\uparrow$3.2for Claude) and Non-RetrICL. We conclude that naively adding ICL examples is not an effective use of the LLM’s context window. Instead, a better content sourcing strategy improves the student distillation, leading to better test performance.
Method Retriever Teacher Self-BLEU-5 ()($\downarrow$)( ↓ ) Entity Entropy ()($\uparrow$)( ↑ ) Mauve ()($\uparrow$)( ↑ ) Accuracy ()($\uparrow$)( ↑ )
(Dataset) LLM AG. IMDb SST-2 AG. IMDb SST-2 AG. IMDb SST-2 AG. IMDb SST-2
Gold - - 17.1 27.9 35.5 6.6 7.5 3.2 - - - 90.8 91.3 88.2
SunGen - GPT2-XL
[…]
15.4
[…]
4.9
[…]
68.7
[…]
84.9
[…]
ReGen BERT - 56.5
[…]
8.1
[…]
68.1
[…]
82.7
[…]
S3 - GPT3.5
$\otimes$
62.2
$\otimes$
$\otimes$
5.7
$\otimes$
$\otimes$
62.0
$\otimes$
$\otimes$
87.1
$\otimes$
AttPmt - GPT3.5-T 39.8
[…]
71.5 6.0
[…]
3.4 52.8
[…]
50.0 79.8
[…]
80.8
Zero-shot
SynzthRR Contr. LLaMa2 29.3 66.3 41.9 7.1 5.7 4.5 89.5 58.5 50.0 85.3 82.9 80.2
SynzthRR Contr. ClaudeV1 31.5 51.5 45.3 6.6 5.3 4.8 94.2 55.9 50.0 85.6 83.6 82.5
SynzthRR BM25 LLaMa2 28.7 62.2 36.5 7.0 5.6 5.1 90.3 60.5 50.0 84.3 74.1 84.4
SynzthRR BM25 ClaudeV1 30.9 50.4 36.9 6.5 5.1 5.4 90.8 53.2 50.0 84.2 79.1 82.6
3-shot RetrICL
SynzthRR Contr. LLaMa2 34.2 62.9 26.3 7.2 5.7 3.8 92.6 72.6 50.0 84.6 84.8 83.8
SynzthRR Contr. ClaudeV1 23.7 38.0 24.6 6.7 5.9 4.3 95.8 58.0 50.0 86.0 86.3 80.6
SynzthRR BM25 LLaMa2 32.0 59.7 25.3 7.2 5.6 4.8 92.5 78.7 50.0 84.3 84.7 84.4
SynzthRR BM25 ClaudeV1 24.6 41.9 26.8 6.7 5.4 4.9 96.0 58.5 50.0 84.1 81.6 82.3
Table 7: Evaluations of synthetic datasets released by prior work. We subsample all to 6K examples (uniformly distributed across classes) before computing metrics as described in §4. Tasks not evaluated by previous authors are denoted by
$\otimes$
while those evaluated without dataset release are marked
[…]
. GPT3.5 is text-davinci-003 whereas GPT3.5-T is gpt-3.5-turbo (OpenAI, 2022), LLaMa2 is 13B Chat version (Touvron et al., 2023a), ClaudeV1 is Instant-V1.2 version (Anthropic, 2023). Accuracy is measured on a DistilBERT student, where we train 5 student models and report the mean accuracy (std. dev. was $\leq 2.0$ in all cases). Within each dataset, we bold the best result.
7 Comparison to previous work
We benchmark SynthesizRR against four prior synthesis methods: (1) SunGen Gao et al. (2023) uses ZeroGen to create 200k synthetic rows and employs a custom bi-level optimization algorithm to weight each instance; (2) ReGen Yu et al. (2023b) utilizes two BERT models, one for retrieval and one as a classifier, to multi-round filter noisy data; (3) S3 Wang et al. (2023a) builds and iteratively enhances a seed dataset by identifying and synthesizing corrections using an LLM; (4) AttrPrompt Yu et al. (2023a) improves dataset diversity and unbiasedness by prompting GPT3.5-Turbo with varied attributes (derived from a human-in-the-loop analysis of each task). Standard zero-shot and few-shot generation baselines were compared in Table 6, so we do not include them here. ZeroGen (Ye et al., 2022a) is similarly excluded.
We benchmark three popular tasks: IMDb Maas et al. (2011), SST-2 Socher et al. (2013) and AG News (Zhang et al., 2015). Previous studies have generated larger datasets ranging from 20k to 200k examples with varying student model hyperparameters, but often lack reports on intrinsic dataset quality, making a fair comparison challenging. Therefore, we independently reproduce these results using the synthetic datasets released by the original authors [^2] ProGen (Ye et al., 2022b) was excluded as it does not release datasets.. Following Yu et al. (2023a), we subsample these datasets to 6k rows, keeping a uniform distribution across classes, and generate the same number of synthetic covariates using SynthesizRR RetrICL (Algorithm 1). For the content sourcing stage of SynthesizRR, we retrieve documents from the CMU Movie Summary corpus Bamman et al. (2013) and RealNews/Dom(Appendix I). We measure accuracy on a DistilBERT student (Sanh et al., 2019; Yu et al., 2023a; Ye et al., 2022a; Gao et al., 2023; Wang et al., 2023a; Ye et al., 2022b), fixing hyperparams to Yu et al. (2023a).
RQ: How does SynthesizRR perform against prior methods on student model accuracy?
Methods like SunGen rely on relatively weak LLM teachers like GPT2-XL Radford et al. (2019) can perform well on topic and sentiment tasks like IMDb, but require a very high data cost (15-30x more synthetic data than SynthesizRR). In Table 7, we observe that when scaled down to 6k rows, the performance deteriorates significantly. We hypothesize that adding the student model into the synthesis process impacts the final classification accuracy, as the dataset becomes specialized to the particular choice of student and does not generalize to other students.
Approaches which use strong instruction-following LLMs like AttrPrompt, S3, and SynthesizRR can achieve similar or better performance with much smaller datasets, as they create high-quality datasets. Prompting techniques like Chain-of-Thought (Wei et al., 2022) used by S3 further improve the task-inversion step (while necessitating higher API costs due to longer output lengths). Chain-of-Thought prompting thus seems like a promising approach to augment SynthesizRR’s task-inversion step.
RQ: do we find evidence that content sourcing promotes diversity and similarity?
Table 7 compares diversity (Entity-Entropy, Self-BLEU), and similarity to Gold texts (MAUVE). Only AttrPrompt (Yu et al., 2023a, Appendix E) attempts to improve diversity of the generated text, by templatizing the task inversion instruction with attributes such as style, topic, length:min-words and more. ReGen is the only prior approach to use content sourcing (but not task inversion). These are thus the most relevant baselines for SynthesizRR.
Both ReGen and SynthesizRR achieve very high entity entropy compared to AttrPrompt, underscoring the importance of a content sourcing step. Unlike SynthesizRR, ReGen uses only retrieval without task-inversion, and thus suffers in terms of lexical diversity, MAUVE and student accuracy.
On the other hand, CoT-style prompting (S3) suffers a lack of lexical diversity and similarity to Gold texts, despite strong distillation performance. This is reproduced in AttrPrompt and previously in FewGen, lending evidence to our claim that synthesis without content sourcing tends to produce datasets with lower diversity, which cannot be overcome by complex prompting strategies alone.
Finally, SunGen exhibits high diversity on IMDb, a task for generating sentiment-based movie reviews. Unlike traditional zero-shot generation, SunGen begins by creating a movie with the prompt Movie: followed by generating an example using prompt The movie review in positive sentiment for movie "<Movie>" is: (details in Ye et al. (2022a, Section 4.6)). We posit that this generated movie fulfils a similar purpose to a retrieved context, enhancing the diversity.
8 Related Work
Dataset synthesis using LLMs. Using LLMs to perform task inversion for dataset synthesis has been studied previously. Most use GPT-2XL without fine-tuning (Ye et al., 2022b, a; Gao et al., 2023; Meng et al., 2022; Schick and Schütze, 2021; Jung et al., 2023). Recent work has considered large teacher LLMs such as GPT-3 (West et al., 2022; Honovich et al., 2023; Wang et al., 2023b), PaLM-540B (Hsieh et al., 2023) and chat-tuned LLMs such as gpt-3.5-turbo (Yu et al., 2023a; Yehudai et al., 2024b; Wang et al., 2023a).
For the generation of text classification datasets, class-conditioned prompting is key. Prior approaches investigated zero-shot (Ye et al., 2022a) and iterative few-shot prompting (Ye et al., 2022b), or synthesis using seq2seq LLMs fine-tuned on a curated dataset (Lee et al., 2021). Recently, AttrPrompt (Yu et al., 2023a) established that varying prompt attributes improves diversity. Our work explores adding retrieval contexts as the source of diversity.
Retrieval-augmented generation. Our approach has many of the characteristics of in-context retrieval-augmented generation (RAG) (Lewis et al., 2020; Ram et al., 2023; Huang et al., 2023; Izacard et al., 2023). Previous studies show how RAG bypasses numerous problems associated with generating solely from parametric memory, i.e., heightened bias towards “head” entities (Mallen et al., 2023), lower lexical diversity (Holtzman et al., 2019; Jentzsch and Kersting, 2023), and hallucinated information (Zhang et al., 2023).
Using retrieval-augmented generation for synthesis of classification tasks has not been explored at the instance level. ReGen (Yu et al., 2023b) studies the retrieval-only setting for creation of topic and sentiment datasets, which are simpler than the tasks in our work. Viswanathan et al. (2023) and Gandhi et al. (2024) perform dataset-level retrieval and not instance-level retrieval.
9 Conclusion
In this work we describe how a retrieval corpus can be used to aid the synthesis of a text classification data set in specialized domains. We show that the diversity of the generated data is enhanced by including retrieved documents in a generation prompt. Compared to few-shot generation, we find that SynthesizRR produces more diverse and representative text and leads to better students.
Limitations
Most principally, our work relies on the existence of a large corpus that is close enough to the task at hand. This may be prohibitive for doing dataset generation in low-resource languages, where a large corpus of related content may not be available. It would be intriguing to explore cross-lingual transfer of content sourcing, but this would require additional experimental validation. By contrast, approaches like FewGen do not require this corpus.
The need for an explicit context sourcing step and increased prompt-length causes an increase in the expenses and latency, especially when using LLM APIs. Such increased expense may not be worth it in the presence of a poor quality retrieval corpus. For one, if the in-context examples are not easily reusable as queries, then SynthesizRR can retrieve irrelevant documents which might not be suitable for task inversion. Furthermore, in the case of factually dubious corpus documents, the student model may end up grounding in factually incorrect information. This can be mitigated by a human-in-the-loop step to remove such documents before task inversion.
Finally, we note that the scope of our experiments is restricted to a set of classification tasks over a few English domains of text. While we believe our approach can be applied to other languages, other domains, and tasks like question answering that go beyond classification, we have not validated this in this work.
Appendix A Risks
Although the main goal of our work is to improve text classification, our use of LLMs to generate examples does carry some conceptual risks. By generating news articles to train classifiers on, we run the risk of generating fake news and other harmful content. However, we believe this risk is mitigated by the fact that the final outcome of our system is a classifier: classification models have relatively constrained failure modes (misclassification) compared to text generation models that can mislead users. Furthermore, we do not believe our approach uniquely advances the generation of content like fake news; our advances are largely orthogonal to the technology that brings such risks.
Appendix B Incorporating feedback from distilled student models
RQ: Why does SynthesizRR improve classification dataset synthesis? In this section we take a closer look at the generated classification dataset and how it affects the training dynamics of student models during distillation.
Aside from the final accuracy, we also consider label preservation accuracy, which is obtained from an “oracle” model for the task. We construct this oracle from Gold data by running a grid-search over DeBERTa-v3-Large hyperparams (Appendix J), splitting 80% of the Gold train set for fine-tuning and 20% for validation. Then, we measure the fraction of synthetic examples which the oracle classifies to belong to the prompted target class. This indicates the adherence of the generated example to the class it should belong to, as per the prompt.
We would expect that better label preservation means a higher-fidelity training dataset. However, Table 8 shows that FewGen datasets have very high label preservation in spite of their lower test performance. Especially on multiclass tasks (AG., ToI, Cat.), FewGen shows the highest label preservation (exceeding Gold) but this does not translate into improved student performance.
Figure 5: Data maps from a DistilBERT training run on $8$K Category rows from LLaMa2. FewGen (center) is skewed towards easy-to-learn examples (top-left) while Gold (left) and SynthesizRR (right) have a higher density of ambiguous examples.
To understand this, we conduct a deeper analysis of the student training dynamics on multiclass datasets. We train a DistilBERT student for 6 epochs and plot the corresponding data-maps Swayamdipta et al. (2020). For binary tasks, the data-maps for SynthesizRR matched both FewGen and Gold, but the data maps from multi-class differed greatly. Figure 5 illustrates this difference using the Category task maps. From Figure 5 it is clear that FewGen generations tend to cluster around easy-to-learn examples (high confidence and low variability), whereas SynthesizRR contains more ambiguous examples (high variability) which Swayamdipta et al. (2020) demonstrate is essential to learning the nuances between classes.
RQ: Can we improve distillation performance by leveraging student feedback from data-maps?
Swayamdipta et al. (2020) use data-maps to filter out easy to-learn examples (top-left, red) and potentially mislabelled examples (bottom-left, blue) and obtain superior accuracy on human-generated datasets. We attempt to apply this same technique to the synthetic datasets generated by SynthesizRR and FewGen.
Concretely, we filter out the least ambiguous examples (bottom 17% variability) and retrain the DistilBERT student model on the smaller, filtered dataset. In Table 9 we find that FewGen performance degrades, whereas SynthesizRR improves (giving us new best performances on multi-class despite using only 83% of rows). We conclude that SynthesizRR generates more ambiguous examples, and this helps establish better class-separability in multi-class data sets.
Method AG. Hyp. ToI Cat. Hum. Pol.
(Dataset size) ($8$K) ($2$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold $93.81$ $81.55$ $85.23$ $84.8$ $95.5$ $96.6$
LLaMa2 Few shot
FewGen* 92.4 $71.25$ 85.9 88.1 $71.7$ $94.75$
SynzthRR $86.88$ 78.6 $74.29$ $72.11$ $90.65$ $94.77$
SynzthRR $87.64$ $75.5$ $74.85$ $74.46$ 95.7 97.6
ClaudeV1 Few shot
FewGen* 94.5 $63.75$ 87.4 89.4 $85.85$ $99.55$
SynzthRR $87.56$ 72.8 $74.79$ $69.41$ 90.7 $99.32$
SynzthRR $87.36$ $65.85$ $73.24$ $73.24$ $77.35$ 99.7
Table 8: Few-shot label-preservation accuracy ($\uparrow$) using tuned oracle DeBERTa-v3L model. Gold row is accuracy on 20% validation split. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
Method AG. ToI Cat. Avg
(Dataset size) ($6.6$K) ($6.6$K) ($6.6$K)
LLaMa2 Few shot
[…]
ClaudeV1 Few shot
[…]
Table 9: Test Accuracy ()($\uparrow$)( ↑ ) after keeping 83% most-ambiguous examples. We report improvements compared to Table 6. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
Appendix C Bootstrapping with a synthetic seed set
A core assumption in SynthesizRR has been the existence of a small seed set of human-written $(x,y)$ pairs for the task. This seed set is critical as it serves a dual purpose: it is used as the set of the retrieval queries, and as in-context learning examples to guide the teacher LLM’s next-token distribution in the task inversion step.
In this section we consider how we can synthesize such a seed set for low-resource settings. Our core assumption is that the seed set is small (100/class for binary tasks and 50/class for multiclass tasks). Thus using FewGen with top-$p=0.9$ and temperature $=0.95$ and three in-context examples, we attempt to generate a diverse seed set with minimal repetitions. This bootstrapping approach makes SynthesizRR tractable when very little human data is available (just 5-15 examples per class) or no human data is available.
Concretely, we compare three paradigms:
  1. 1.
    True zero-shot: when we have no human data we utilize zero-shot generation to bootstrap the seed set.
  2. 2.
    Low-resource: Here, we assume we have a small number of human-written examples, e.g. 5 examples per class. This is presumed insufficient to be used as the seed set directly, but we can use it as in-context examples to guide the FewGen generator to bootstrap a realistic seed set.
  3. 3.
    Sufficient: We do not synthesize the seed set. This is the SynthesizRR paradigm we have explored in previous sections, wherein we have 50-100 Gold examples per class in our seed set.
As mentioned in §4, the true zero-shot paradigm makes strong assumptions that are often unnecessarily restrictive. In practice, it is typically feasible to obtain a small amount of human-written examples (low-resource or sufficient seed), while obtaining several thousand human-written examples is still challenging.
Gold RetrICL AG. Hyp. ToI Cat. Hum. Pol.
data ($N$) shots ($8$K) ($2$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold
All - 91.0 93.2 82.5 81.5 93.1 95.3
True Zero-shot (0-shot FewGen seed)
None 0-shot 66.6 68.0 60.5 60.4 76.9 76.4
None 3-shot 60.0 72.3 62.5 61.7 72.3 85.4
Low-Resource ($\binom{N}{3}$-shot FewGen seed)
5/class 0-shot 79.9 71.7 68.1 63.4 81.3 81.3
5/class 3-shot 77.7 66.8 68.9 58.8 86.4 86.5
15/class 0-shot 78.5 72.9 69.3 65.7 77.4 84.0
15/class 3-shot 76.1 72.6 71.6 63.5 82.5 73.8
Sufficient (Gold Seed)
Full seed 0-shot 83.5 69.8 74.5 68.9 82.5 84.7
Full seed 3-shot 83.0 78.5 73.3 72.4 90.2 91.0
Table 10: Test accuracy after distilling a DeBERTa-v3L student on a dataset generated by SynthesizRR RetrICL variant. We use the same corpus as Table 2, but vary the seed set. LLaMa-2 Chat 13B is used as the teacher LLM. We train 5 student models and report the mean accuracy, rerunning all 5 in case of std $\geq$ 6.0. “’Full” seed implies 100 Gold examples per class for binary and 50 per class for multiclass tasks. We bold the best result in each paradigm.
The results of running SynthesizRR RetrICL using synthetic seed data is shown in Table 10. As a general trend, adding more human-written examples leads to better performance. Unsurprisingly, the best results are in the Sufficient paradigm, where we use 50-100 Gold examples as both retrieval queries and the the RetrICL set. True Zero-shot results (without any human input) are considerably worse. Surprisingly, however, we are able to get good distillation accuracy with just 5 examples per class rather than the full 50-100 per class, which indicates that SynthesizRR might be usable in low-resource settings where human annotated data is scarce.
In certain cases of the low-resource paradigm, we observe that the performance drops significantly from 0-shot RetrICL to 3-shot RetrICL. We attribute this to the fact that, even with 5-15 Gold in-context examples, the FewGen-generated seed set might not be reflective of the true Gold examples (this behavior is reflected in the low MAUVE scores in Table 5). Thus, by conditioning on incorrect synthetic examples during RetrICL, we shift the next-token distribution away from the true distribution.
In conclusion, using FewGen to bootstrap a seed set can be a viable approach to using SynthesizRR in low-resource settings where there is not enough Gold task-data.
Appendix D Influence of corpus on domain shift
AG News ($4$K)
Corpus DeBERTa ($\uparrow$) Mauve ($\uparrow$) Self-BLEU-5 ($\downarrow$) Entity Ent. ($\uparrow$)
RN/Dom 85.39 ± 0.8 92.58 0.23 6.72
RN/Rnd 35.57 ± 6.1 83.39 0.22 7.07
RN/Reg 84.17 ± 0.7 88.88 0.26 6.72
Hyperpartisan ($2$K)
Corpus DeBERTa ($\uparrow$) Mauve ($\uparrow$) Self-BLEU-5 ($\downarrow$) Entity Ent. ($\uparrow$)
RN/Dom 78.77 ± 2.8 66.94 0.35 6.11
RN/Rnd 78.77 ± 3.5 61.45 0.25 7.40
RN/Reg 72.00 ± 2.0 65.59 0.35 6.12
Table 11: Effect of corpus-swapping for SynthesizRR 32-shot Non-RetrICL. We generate only $4$k rows for AG News to reduce costs.
Our expectation is that SynthesizRR can flexibly specialize students to different domains by transparently changing the retrieval corpus, while keeping a frozen LLM. To quantify how changing the retrieval corpus might affect earlier metrics, we switch the news corpus for Hyperpartisan and AG News. We had assumed RealNews/Dom was the most suitable corpus (in-domain), and the others will cause domain-shift. In the following RQs, we validate the degree to which this assumption holds and the importance of information retrieval as the content sourcing mechanism in SynthesizRR.
RQ: Does modifying the corpus cause domain shift? Table 11 finds that the retrieval corpus highly influences the test performance (both student and intrinsic metrics). When grounding to a corpus with highly dissimilar entities (such as RealNews/Reg), all metrics drop significantly. Thus, we can conclude that an alternative content-source does indeed induce domain-shift. Mauve and distillation accuracy are highest for the in-domain corpus, while Self-BLEU and Entity entropy are highest for the random-retrieval results.
Figure 6: Retrieval counts for Hyperpartisan and AG News. The red dashed line represents the theoretical max, where all retrieved documents are unique. Note that the Random histogram plot is always 1 hence shows up as a straight line.
RQ: is retrieval essential for content sourcing? We measure the importance of retrieval by selecting top-k documents randomly from the in-domain corpus RealNews/Dom. We observe in Table 11 that retrieval using in-context learning queries plays a crucial role to the performance of AG News, as performance drops significantly in a random setting. Hyperpartisandoes not face such a drop. This matches our intuition in Table 1 that task-inversion is the more challenging step for Hyperpartisan, and a powerful LLM we can apply stylistic changes to most news articles. In both, Mauve suffers when entities no longer match Gold.
RQ: Do in-context queries retrieve redundant results? Figure 6 measures the overlap of top-50 retrieved documents from the 200 ICL queries, and finds that in most cases, the retrieved documents are unique, especially when using a large in-domain corpus. Thus, we can conclude that effective retrieval is important for the diversity of the synthetic dataset.
RQ: Can SynthesizRR work effectively with relatively small corpora? In our main results §5, we assumed the existence of a large corpus, with minimum size of 0.9M documents. As noted, this corpus need not be unlabelled examples for our task; we were able to successfully generate customer reviews and product questions for Humor, Category and Polarity tasks, while retrieving from a corpus of product information (title and description).
A potential problem with SynthesizRR is that corpuses of such massive size might be few in number. Thus, we compare the performance of SynthesizRR on CMU Movie Summary Bamman et al. (2013) which is between one to three orders of magnitude smaller than other corpora in Table 6. In Table 7, we see that SynthesizRR can perform suitably even with such relatively small corpora (42k movie plots). From the previous RQs, this suggests that the relevance of the corpus to the task is more important than the size of the corpus for the performance of SynthesizRR.
Appendix E Dense vs sparse retrieval in SynthesizRR
Retriever AG. Hyp. ToI Cat. Hum. Pol. Avg.
(Size) ($8$K) ($2$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold 91.0 93.2 82.5 81.5 93.1 95.3 89.43
LLaMa2 Zero shot
Contr. 83.5 69.8 74.5 68.9 82.5 84.7 77.32
BM25 83.2 74.2 70.7 57.6 78.5 85.4 74.93
ClaudeV1 Zero shot
Contr. 83.9 72.3 71.8 66.8 62.1 88.7 74.29
BM25 83.2 57.2 69.8 53.7 73.9 91.8 71.60
LLaMa2 3-shot RetrICL
Contr. 83.0 78.5 73.3 72.4 90.2 91.0 81.38
BM25 82.1 77.9 71.9 65.4 87.5 87.4 78.69
ClaudeV1 3-shot RetrICL
Contr. 83.7 72.3 72.8 65.4 83.4 91.3 78.16
BM25 83.0 73.5 70.0 52.4 82.4 90.7 75.34
Table 12: Test accuracy after distilling a DeBERTa-v3L student on a dataset generated by SynthesizRR. Retrieval is done using BM25 and Contriever. We use the same seed set and corpus as Table 2. We train 5 student models and report the mean accuracy, rerunning all 5 in case of std $\geq$ 6.0. The top two subsections consider zero-shot synthesis and bottom two considers 3-shot RetrICL variant. We bold the best result in each subsection. Contriever numbers are reproduced from Table 6.
So far, a single dense retriever (Contriever) has been used for the content sourcing step by using a bi-encoder approach (Lee et al., 2019; Chen et al., 2017). We embed both the input in-context covariate and each corpus document, and then rank results based on cosine similarity. In §5, we retrieved $k=500$ documents for each in-context example and after filtering, randomly sampled among these to produce a grounded set of documents on which we apply our task inversion strategy RetrICL.
In this section we explore how changing the retrieval model affects the content sourcing stage and its downstream effects. Keeping other parts of the process the same, we switch Contriever to BM25 Okapi (Robertson and Zaragoza, 2009), a popular sparse retrieval method. Dense retrievers like Contriever perform a semantic match between the query and document, whereas BM25 performs only a lexical match based on inverse term frequencies, with no understanding of semantics. Additionally, BM25 outputs a score which is an unbounded positive number, thus we are unable to use meaningful thresholds to bound the similarity in our RetrICL approach. Instead, we construct the RetrICL in-context set using the top-2 retrieved contexts for each ICL example and without applying the filter.
We expect that picking semantically similar information is more important to SynthesizRR since we include a task inversion step, which intends to change the tone and lexical structure of the text while preserving its semantics. Thus, we want contexts which are semantically related to Gold data, to which we can apply stylistic or formatting transformations using a task-inversion prompt to bring it closer to Gold.
Surprisingly, in Table 7 we see that while intrinsic diversity from BM25-retrieved documents is often worse than Contriever, they both generate equally human-like text. However, comparing the DeBERTa-v3L accuracy of Contriever and BM25in Table 12, we see that a strong student model trained on a dataset obtained from the dense-retrieved document set consistently outperforms the sparse retriever BM25, which might be due to the filtering step we introduce in RetrICL. This filtering step leads to a reduction in mislabelling stemming from retrieving contexts that belong do a different class. Due to this, we conclude that dense retrieval models are potentially more suitable for SynthesizRR.
Appendix F Varying number of in-context examples in RetrICL
Figure 7: Left: DeBERTa-v3L test accuracy ()($\uparrow$)( ↑ ), center: entity entropy ()($\uparrow$)( ↑ ), right: Mauve ()($\uparrow$)( ↑ ) for SynthesizRR RetrICL. We vary the number of in-context examples from 0 to 8. Teacher LLMs LLaMa-2 Chat 13B and Claude Instant-v1 are compared on 6 tasks: AG News, Hyperpartisan, ToI Headlines, Category, Humor and Polarity. We do not report Category 8-shot due to model failures.
[…]
The use of in-context examples in the RetrICL variant of SynthesizRR leads to significant improvements in intrinsic and distillation metrics, as we saw in §5. Here, we do a deeper analysis on whether continually increasing the number of in-context examples yields a positive benefit.
In Figure 7 we look at the DeBERTa-v3L accuracy, entity entropy and MAUVE for our datasets with different numbers of in-context learning examples. We see that adding even a single in-context example can greatly increase the performance of all three metrics. However, no particular number of in-context examples consistently outperforms. For ClaudeV1, adding more in-context examples (up to 8) seems to always provide benefit, whereas with LLaMa2, we observe a peak and then reduction. Thus, the optimal number of in-context learning examples is a task dependent hyperparameter.
Figure 8 shows the lexical diversity i.e. Self-BLEU across datasets and number of in-context examples. As in §5 we observed that using in-context examples is neither positively nor negatively correlated with a lower Self-BLEU, despite using nucleus sampling with $p=0.9$. This may be because for all number of shots, task inversion is performed from a single source context and thus the generation does not divert significantly from the unique n-grams of the context. Thus we conclude that to affect lexical diversity, the number of in-context learning examples has no effect and we must instead focus on changing the retrieved contexts, perhaps by using a different retrieval model.
Appendix G Task inversion prompts and label verbalizations
Here we discuss the prompt templates and verbalizations that we use for the task inversion step for both FewGen and SynthesizRR. We use descriptive verbalizations as compared to the target label.
Additionally in the prompt, we place the retrieved document near the end, as prior work indicates that intermediate placements degrade LLM recall (Liu et al., 2023).
LLMs have a fixed window-size for conditional generation, so excessively long documents are truncated (from the end) up to $r_{max}=500$ tokens. This reserves the remaining window for in-context learning.
G.1 Hyperpartisan
Hyperpartisan is the task of detecting political bias in a news article. In transforming the retrieved news article article_retr[k] to one with such bias, typically there is the addition of mocking commentary and harsh political language which deeply criticizes the subject such as a person, policy or political event. On the other hand, articles in the opposite class gives a well-rounded opinion with a neutral tone. We include a length-attribute to ensure a long generation of one or two paragraphs.
Label Verbalization
true harsh political language, using a mocking tone and toxic commentary
false neutral language, using a reasonable tone and politically correct commentary
Table 13: Task-inversion verbalizations for Hyperpartisan.
Prompt G.1: Hyperpartisan FewGen In-context example:
Write a single news article using {label} . The written article should be 2 to 3 paragraphs long.
News Article: {icl[gold_text]}
Prompt:
Write a single news article using {label} . The written article should be 2 to 3 paragraphs long.
News Article:
Prompt G.2: Hyperpartisan SynthesizRR RetrICL In-context example:
News Article: {icl[article_retr]}
Rewrite the above news article using {label} . The rewritten article should be 2 to 3 paragraphs long.
Rewritten Article: {icl[gold_text]}
Prompt:
News Article: {article_retr[k]}
Rewrite the above news article using {label} . The rewritten article should be 2 to 3 paragraphs long.
Rewritten Article:
Prompt G.3: Hyperpartisan SynthesizRR Non-RetrICL In-context example:
Rewritten Article: {icl[gold_text]}
Prompt:
News Article: {article_retr[k]}
Rewrite the above news article using {label} . The rewritten article should be 2 to 3 paragraphs long.
Rewritten Article:
G.2 ToI Headlines
ToI Headlines is a topic classification dataset of regional news headlines in India. Here we attempt to refine the retrieved news article by summarizing it into a short headline. We use verbalizations of the content of each class, as example generation here involves summarizing the content. We add an “India” location-attribute to guide the LLM generations to include regionalization to the Indian subcontinent. A length-attribute is included to restrict the length to one sentence.
Label Verbalization
sports sports in India
life-style health and lifestyle trends in India
education Indian examinations and education
entertainment the Indian entertainment industry
business business-related developments in India
city ongoing matters in any Indian city
environment environment-related events in Indian cities
tech technology news and the tech industry in India
elections elections and politics in India
world international news and events outside of India
Table 14: Task-inversion verbalizations for ToI Headlines.
Prompt G.4: ToI Headlines FewGen In-context example:
Write a headline for a news article about {label} . The headline should be a single sentence.
Headline: {icl[gold_text]}
Prompt:
Write a headline for a news article about {label} . The headline should be a single sentence.
Headline:
Prompt G.5: ToI Headlines SynthesizRR RetrICL In-context example:
News Article: {icl[article_retr]}
Write a headline for the above news article about {label} . The headline should be a single sentence.
Headline: {icl[gold_text]}
Prompt:
News Article: {article_retr[k]}
Write a headline for the above news article about {label} . The headline should be a single sentence.
Headline:
Prompt G.6: ToI Headlines SynthesizRR Non-RetrICL In-context example:
Headline: {icl[article_retr]}
Prompt:
News Article: {article_retr[k]}
Write a headline for the above news article about {label} . The headline should be a single sentence.
Headline:
G.3 AG News
We consider task inversion for the AG News dataset to be generation of news summaries. We do not specify location modifiers as most Gold summaries are from US news. We add a length-attribute to restrict the output one or two sentences.
Label Verbalization
Business companies, industries, markets, trade, investments, entrepreneurship, economic policies, and other business-related developments
World international news, such as politics, diplomacy, conflicts, global events, international relations, human rights issues, and significant global trends
Sci/Tech scientific discoveries, technological advancements, innovations, research breakthroughs
Sports professional sports leagues, major tournaments, athletes, teams, match results, player transfers, coaching changes, sports-related controversies
Table 15: Task-inversion verbalizations for AG News.
Prompt G.7: AG News FewGen In-context example:
Write a summary for a news article about {label} . The summary should be one or two short sentences.
Summary: {icl[gold_text]}
Prompt:
Write a summary for a news article about {label} . The summary should be one or two short sentences.
Summary:
Prompt G.8: AG News SynthesizRR RetrICL In-context example:
News Article: {icl[article_retr]}
Write a summary for the above news article about {label} . The summary should be one or two short sentences.
Summary: {icl[gold_text]}
Prompt:
News Article: {article_retr[k]}
Write a summary for the above news article about {label} . The summary should be one or two short sentences.
Summary:
Prompt G.9: AG News SynthesizRR Non-RetrICL In-context example:
Summary: {icl[gold_text]}
Prompt:
News Article: {article_retr[k]}
Write a summary for the above news article about {label} . The summary should be one or two short sentences.
Summary:
G.4 Category
In the Category dataset, we determine the product category from a review written by a user about products on a major e-commerce website. For task inversion in SynthesizRR we must retrieve a product and prompt the frozen LLM to generate a user review within the same product-category as the retrieval query. Thus, we include a style-attribute to allow minor typos in the generation and restrict to a few sentences using a length-attribute.
Label Verbalization
magazines magazines or periodicals covering various topics
camera_photo photography gear including cameras, lenses, accessories, or photo editing tools
office_products office supplies or equipment for professional and home office setups
kitchen kitchenware, appliances, or culinary tools for cooking and dining
cell_phones_service cell phone service accessories or service plans for communication and connectivity
computer_video_games computers, gaming consoles, video games, or related accessories
grocery_and_gourmet_food groceries, fruits and vegetables, gourmet treats, or specialty food items
tools_hardware tools, hardware, or equipment for DIY projects and home repairs
automotive auto parts, accessories, or tools for vehicle maintenance and enhancements
music_album music albums spanning various genres and artists
health_and_personal_care healthcare products, personal care items, or wellness essentials
electronics electronic devices, gadgets, personal tech, or home electronics
outdoor_living products for outdoor activities, gardening, or patio living
video movies, TV shows, and documentaries spanning various genres and artists
apparel clothing including casual wear, formal attire, seasonal outfits, activewear, or fashion accessories for men, women, and children
toys_games fun or educational toys and games for kids of all ages
sports_outdoors products for various sports and outdoor activities
books books in various genres and formats
software computer software for productivity or gaming covering either personal or professional needs
baby baby essentials, gear, or toys for infants and toddlers
musical_and_instruments musical instruments, accessories, or music production equipment
beauty beauty products, cosmetics, or skincare essentials, makeup, hair care, fragrances, or grooming essentials
jewelry_and_watches watches or jewelry pieces such as necklaces, bracelets, earrings, or rings, crafted in precious metals or adorned with gemstones for special occasions
Table 16: Task-inversion verbalizations for Category.
Prompt G.10: Category FewGen In-context example:
Write a product review about a product which is in the category of {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Write a product review about a product which is in the category of {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.11: Category SynthesizRR RetrICL In-context example:
Product details: {icl[product_retr]}
Write a product review about the above product which is in the category of {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Product details: {product_retr[k]}
Write a product review about the above product which is in the category of {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.12: Category SynthesizRR Non-RetrICL In-context example:
Review: {icl[gold_text]}
Prompt:
Product details:
Write a product review about the above product which is in the category of {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
G.5 Humor
Asking humorous product questions is a challenge of the LLM’s task inversion capabilities, as it must generate a question which is funny from the retrieved product. Not all products have obvious humorous characteristics, thus the generation requires some ingenuity. We restrict the output to only the question to prevent explanations or extraneous product generations from the LLM.
Label Verbalization
humorous humorous
non_humorous solemn
Table 17: Task inversion verbalizations for Humor.
Prompt G.13: Humor FewGen In-context example:
Write a short {label} question about a product. Only include the question.
Product Question: {icl[gold_text]}
Prompt:
Write a short {label} question about a product. Only include the question.
Product Question:
Prompt G.14: Humor SynthesizRR RetrICL In-context example:
Product details: {icl[product_retr]}
Write a short {label} question about the above product. Only include the question.
Product Question: {icl[gold_text]}
Prompt:
Product details: {product_retr[k]}
Write a short {label} question about the above product. Only include the question.
Product Question:
Prompt G.15: Humor SynthesizRR Non-RetrICL In-context example:
Product Question: {icl[gold_text]}
Prompt:
Product details: {product_retr[k]}
Write a short {label} question about the above product. Only include the question.
Product Question:
G.6 Polarity
Polarity is a sentiment classification task for reviews of products on a major e-commerce website. In SynthesizRR, the difficulty is increased as we must generate a review from a product. For task inversion, we prompt the LLM to generate a review which can have either positive or negative sentiment and include details from the retrieved product. As with Category, we allow typos and restrict the length to a few sentences using a length-attribute in the prompt.
Label Verbalization
positive what the reviewer liked about the product, how the reviewer found it easy to use the product, or the reviewer’s positive experience with the product
negative what the reviewer disliked about the product, how the reviewer found it challenging to use the product, or the reviewer’s negative experience with the product
Table 18: Task inversion verbalizations for Polarity.
Prompt G.16: Polarity FewGen In-context example:
Write a review about a product which discusses {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Write a review about a product which discusses {label} . Include relevant product details. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.17: Polarity SynthesizRR RetrICL In-context example:
Product details: {icl[product_retr]}
Write a review about the above product which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Product details: {product_retr[k]}
Write a review about the above product which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.18: Polarity SynthesizRR Non-RetrICL In-context example:
Review: {icl[gold_text]}
Prompt:
Product details: {product_retr[k]}
Write a review about the above product which discusses {label} . Include relevant product details which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
G.7 IMDb
IMDb is a review-sentiment classification task. As with other review tasks, in the task inversion step we prompt the LLM to generate a review in either positive or negative sentiment. The context used by SynthesizRR is the plotline of a movie from CMU Movie Summary. As with Category and Polarity, we allow typos and restrict the length to a few sentences using a length-attribute in the prompt.
Label Verbalization
positive what the reviewer liked about the movie
negative what the reviewer disliked about the movie
Table 19: Task inversion verbalizations for IMDb.
Prompt G.19: IMDb FewGen In-context example:
Write a review which discusses {label} . Include relevant details about the movie. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Write a review which discusses {label} . Include relevant details about the movie. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.20: IMDb SynthesizRR RetrICL In-context example:
Movie details: {icl[plotline_retr]}
Write a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Movie details: {plotline_retr[k]}
Write a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
Prompt G.21: IMDb SynthesizRR Non-RetrICL In-context example:
Review: {icl[gold_text]}
Prompt:
Movie details: {plotline_retr[k]}
Write a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence, or a single paragraph of 3 to 4 sentences. Add very minor typos.
Review:
G.8 SST-2
SST-2 is another review-sentiment classification task, however the examples are partial sentences from movie reviews which were extracted such that they contain the sentiment-heavy phrases. This, during the task inversion we prompt the Teacher LLM to generate a partial review sentence in either positive or negative sentiment. The context used by SynthesizRR is the plotline of a movie from CMU Movie Summary. We allow typos and restrict the length to one sentence using a length-attribute in the prompt.
Label Verbalization
positive what the reviewer liked about the movie
negative what the reviewer disliked about the movie
Table 20: Task inversion verbalizations for SST-2.
Prompt G.22: SST-2 FewGen In-context example:
Write a single sentence from a review which discusses {label} . Include relevant details about the movie. The review should only be a single short sentence. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Write a single sentence from a review which discusses {label} . Include relevant details about the movie. The review should only be a single short sentence. Add very minor typos.
Review:
Prompt G.23: SST-2 SynthesizRR RetrICL In-context example:
Movie details: {icl[plotline_retr]}
Write a single sentence from a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence. Add very minor typos.
Review: {icl[gold_text]}
Prompt:
Movie details: {plotline_retr[k]}
Write a single sentence from a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence. Add very minor typos.
Review:
Prompt G.24: SST-2 SynthesizRR Non-RetrICL In-context example:
Review: {icl[gold_text]}
Prompt:
Movie details: {plotline_retr[k]}
Write a single sentence from a review which discusses {label} . Include relevant details about the movie which are mentioned above. The review should only be a single short sentence. Add very minor typos.
Review:
Appendix H Example generations
Here we showcase examples from the best-performing SynthesizRR approach (3-shot Non-RetrICL using LLaMa-2 Chat 13B) for each of our 6 tasks. For brevity, we do not show the ICL examples, only the retrieved article and generated text.
Class Example
(Retrieved document) News Article: Gov. Andrew Cuomo Wednesdy afternoon staged a press conference to blast federal immigration officials, threatening to sue if agents don’t stop what he described as "reckless and unconstitutional enforcement" practices. Immigration and Customs Enforcement Deputy Director Thomas D. Homan responded hours later with a six-paragraph statement that accused Cuomo of "grandstanding." Cuomo cited a raid last week on a farm in Rome, N.Y. Immigration and Customs Enforcement agents entered John Collins’ property and detained a worker, Marcial DeLeon Aguilar, without presenting a warrant to Collins. That worker was later arrested and is still being detained. "The reckless and unconstitutional practices ICE is deploying in our communities violate everything we believe in New York and are an assault on our democracy," Cuomo said in a prepared statement. "I demand ICE immediately cease and desist this pattern of conduct, and if they fail to do so, I will pursue all available legal recourse and commit to doing everything in my power to protect the rights and safety of all New Yorkers." At about 7:30 p.m., Homan issued this response: "As a native New Yorker who began my law enforcement career as a police officer in the state, I was disappointed to learn about the Governor’s grandstanding today over the issue of immigration enforcement. The Governor’s comments were inaccurate and an insult to ICE’s sworn law enforcement officers who conduct their lawful mission professionally and with integrity. These brave men and woman leave the safety of their homes every day to protect this great nation and our communities.
hyperpartisan What a joke! Governor Andrew Cuomo had the audacity to call out ICE agents for doing their jobs, and now he’s threatening to sue them? Give me a break! The nerve of this guy, talking about "reckless and unconstitutional enforcement" practices when his own state is infested with illegal immigrants who are committing crimes left and right. Cuomo is just trying to grandstand and pander to his liberal base, but the fact is, he’s failing to protect the citizens of New York. ICE agents are just doing their jobs, and they deserve respect for their hard work and dedication. Cuomo should be ashamed of himself for attacking these brave men and women. And let’s not forget, this is the same guy who is protecting child rapists and allowing them to roam free in his state. What a disgrace! Cuomo needs to focus on the real problems in his state, like the rampant crime and corruption, instead of trying to score political points by attacking ICE agents. It’s pathetic.
(Retrieved document) News Article: This week on "Watters’ World", Jesse Watters caught up with several House Democrats who boycotted President Trump’s inauguration, including Rep. John Lewis, the Atlanta Democrat who called the Republican’s election "not legitimate". Approaching Houston Congresswoman Sheila Jackson Lee, Watters asked why she was absent from the festivities: "I have a statement on record, thank you so very much," Jackson Lee said. WATCH: Watters Tracks Down Rep. John Lewis to Ask About Trump Criticism ’Watters’ World’ EXCLUSIVE: Jesse Visits Scene of Chicago Hate Crime Attack Watters Confronts Man Who Harassed Ivanka Trump on JFK Flight As Watters introduced himself as a Fox News reporter to Rep. John Conyers of Michigan, the Detroiter kept walking toward an open elevator, "no, sir, not right now." "I didn’t boycott it, I just didn’t go to it," Conyers said. Rep. Jerrold Nadler of New York gave a similar answer when asked, "for the reasons I stated publicly; they’re on the record." Asked if he was being divisive, the Manhattan Democrat said that if anyone wanted to unite the country, they would not have voted for Trump. Lewis told Watters he was late and couldn’t speak with him as he walked with a quick gait down a hallway. An aide obstructed Watters’ cameraman as the Congressman continued toward his destination.
non-hyperpartisan In a recent segment on "Watters’ World", Jesse Watters spoke with several members of the House of Representatives about their absence from President Trump’s inauguration. While some members provided brief statements, others declined to comment or did not respond to Watters’ inquiries. Rep. John Lewis, an Atlanta Democrat who had publicly questioned the legitimacy of Trump’s election, was approached by Watters but did not stop to speak with him. Lewis had previously stated that he would not attend the inauguration due to his disagreement with Trump’s presidency. Other members, including Rep. John Conyers of Michigan and Rep. Jerrold Nadler of New York, also declined to comment or did not respond to Watters’ questions. When asked why he did not attend the inauguration, Conyers simply stated that he "didn’t go to it". Nadler, who had publicly stated that he would not attend the inauguration due to his disagreement with Trump’s policies, reiterated that he did not support the President’s election. Watters’ segment sparked a discussion on the perceived divisiveness of the President and the responses of those who did not attend his inauguration. While some argued that the boycott was a form of protest against Trump’s policies and actions, others felt that it was a divisive move that further polarized the country. Regardless of one’s political beliefs, it is important to approach discussions and debates with respect and civility, and to strive for understanding and cooperation in the pursuit of common goals.
Table 21: Generated examples for Hyperpartisan task using SynthesizRR.
Class Example
(Retrieved document) News Article: New Delhi: First reports of exit poll indicate history in West Bengal and regime change in Kerala while it’s too close to call in Tamil Nadu. An exit poll conducted by Star News television channel has projected Congress-led United Progressive Alliance (UPA)’s victory in West Bengal and Kerala. The survey projects Trinamool Congress - Congress alliance wiping out more than three decades old Left Front government in West Bengal. According to the exit polls telecast in the channel on Tuesday says the TMC and Congress will win 181 and 40 respectively while Left Front will be reduced to 62 from 227. In Kerala, the Congress-led United Democratic Front is expected to win 88 while CPM-led Left Democratic Front will win 49. In Tamil Nadu, AIADMK is projected as winning 110 and ruling DMK will have 124 seats in the 234 member assembly. In a close contest, Tarun Gogoi led Congress has a chance to create history in Assam by returning to power for a third consecutive term,post poll surveys predict. The Headlines Today poll predicts 44 of the 126 seats for Congress in Assam,with potential ally AUDF slated to win 14.
elections Exit Polls predict regime change in Kerala, West Bengal; Tight contest in Tamil Nadu
(Retrieved document) News Article: India’s teen sensation P V Sindhu stormed into the third round of the World Badminton Championship with a straight game victory over Olga Golovanova of Russia in women’s singles match here yesterday Copenhagen: India’s teen sensation PV Sindhu stormed into the third round of the World Badminton Championship with a straight game victory over Olga Golovanova of Russia in women’s singles match here yesterday. PV Sindhu The men’s doubles pair of Manu Attri and Sumeeth Reddy B stunned 15th seeded Japanese duo of Hirokatsu Hashimoto and Noriyasu Hirata 21-19 21-19 in 44 minutes to advance to the third round. Sindhu, seeded 11th, took 40 minutes to prevail over her Russian opponent 21-12 21-17 in the second round match at the Ballerup Super Arena here. She will next take on sixth seeded Yeon Ju Bae of Korea. Sindhu won a total of 42 points as compared to 29 by the Russian girl. The world No. 12 from Hyderabad looked a bit rusty to start with and was initially trailing in the opening game. She was playing the catching-up game till 10-10 before Sindhu managed to reel off four consecutive points and surge ahead to 14-10. There was no looking back after that, as Sindhu swiftly sealed the game in her favour with Golovanova earning just two more points. In the second game, the Russian got her act together and opened up a big lead, moving up to 11-6 at the break.
sports 15-year-old PV Sindhu creates history, enters World Badminton Championship 3rd round
Table 22: Generated examples for ToI Headlines task using SynthesizRR.
Class Example
(Retrieved document) News Article: SANTIAGO Nov 14 A magnitude 6.1 earthquake hit central-north Chile on Wednesday, the U.S. Geological Survey said, causing buildings to sway in capital Santiago but causing no damage. The USGS said the quake struck on the coast, 54 miles (88 km) north of Coquimbo, at 4:02 p.m. (1902 GMT). It was initially reported as a magnitude 6.0 quake. Chile’s state emergency office, Onemi, said there were no reports of injuries, disruptions to basic services or infrastructure damage as a result of the tremor. The navy said the quake did not meet the conditions needed to trigger a tsunami. World No.1 copper producer Codelco said its operations were normal, and no mines in the metal-producing country reported damages. Seismically active Chile was battered in February 2010 by an 8.8 magnitude quake and ensuing tsunami, which caused billions in damage and killed hundreds.
World A magnitude 6.1 earthquake hit central-north Chile on Wednesday, causing buildings to sway in capital Santiago but causing no damage, the USGS said.
(Retrieved document) News Article: MONTREAL - Canadian astronaut David Saint-Jacques is bracing for some tense moments during next Monday’s launch of the Soyuz rocket that will send him and two others to the International Space Station. On Oct. 11, a rocket failure forced a Soyuz capsule carrying two astronauts to abort and make an emergency landing. Russia suspended all manned space launches pending an investigation before giving the green light Nov. 1. Saint Jacques spoke to reporters today from the launch site in Kazakhstan where he’s in quarantine. He said the most dangerous part of the six-month mission is the 10-minute Soyuz launch and the six hours that follow before docking. He said half of the last two-and-half years of training was dedicated to his role as Soyuz co-pilot for the trip to the space station. Saint-Jacques, 48, said once at the space station he’ll be able to focus on work and life aboard the outpost. The Canadian astronaut’s first space voyage, with NASA astronaut Anne McClain and Oleg Kononenkoof of the Russian space agency Roscosmos, was originally scheduled to launch Dec. 20. It was moved ahead after Russian authorities concluded their investigation into the failed launch.
Sci/Tech commentary Saint-Jacques says the Soyuz launch is the most dangerous part of the six-month mission but he’s been training for it for half of the last two-and-a-half years.
Table 23: Generated examples for AG News task using SynthesizRR.
Class Example
(Retrieved document) Product details: MSD ASY17296 Wiring Harness Factory Terminated MSD Ignition 6ALN (6430) Wiring Harness ASY17296 New harness with factory crimped connections a perfect item to have as an extra or replacement.
camera_photo I have a MSD Ignition 6ALN (6430) and this wiring harness is a great addition. The factory terminated connections are perfect and make installation a breeze. The harness is well made and will definitely last. I have been using it for a few weeks now and it has performed flawlessly. Highly recommend it to anyone looking for a high quality wiring harness.
(Retrieved document) Product details: Conair Travel Smart 3-Dial Combination Lock (6 Pack) Safely Locks Your Luggage With Tsa Recognized Locks Comes With A Red Inspection Status Feature Travel Sentry Approved
apparel I purchased this lock for a business trip to Europe. I’ve used it several times already and so far, so good. The TSA inspectors have been able to open it without a hitch, which is great!
Table 24: Generated examples for Category task using SynthesizRR.
Class Example
(Retrieved document) Product details: Keeping Up With The Kardashians - Season 5 Region E! brings viewers back into the sensational and completely unpredictable world of one of televisions favourite families with the fourth season of the hit series Keeping Up with the Kardashians. This season, Khlo and her husband, Los Angeles Lakers star, Lamar Odom, begin their new lives as a married couple. Meanwhile, Kourtney is pregnant and back together with Scott, and Kim and her boyfriend, NFL player Reggie Bush, work on their relationship. Through all this change, the Kardashian family must find ways to evolve and keep their close-knit ties.
humorous Does anyone else notice that this is the only show where you can hear people say things like, "OMG I am totally exhausted, I need to take a nap.? And it’s not considered weird?
(Retrieved document) Product details: Coleman Camper’s Storage Tote Securely organize, store and carry camping gear 23" x 14" x 17" Lightweight 600D polyester construction Multiple pockets and compartments Convenient carry strap The Coleman Camper’s Storage Tote is a great way to store, organize and carry your gear – use it in the garage, in the trunk or RV, and at the campsite. Constructed of reliable, lightweight 600D polyester, the Camper’s Storage measures 23" long, 14" wide and 17" tall. It features multiple pockets and compartments for organizing all your gear, and a convenient carry strap so it’s easy to bring along.
non_humorous is this product big enough for all the camping gear for a family of four?
Table 25: Generated examples for Humor task using SynthesizRR.
Class Example
(Retrieved document) Product details: Birth Of Soul 1 No half measures here, but a real statement of intent, particularly with that Volume One tag. 28 cuts that helped define that thing called soul - you can argue the point, but you can’t argue the music. Sublime.
positive This album was the beginning of my love affair with soul music. These tracks are timeless and still sound amazing today.
(Retrieved document) Product details: Beware! the Lies of Satan It’s high time we pull the covers off the enemy and expose him for who he really is, the murderer, destroyer, their, deceiver and biggest liar of all time. In this book , you will begin to discover the truth about our heavenly Father and how you can stand in victory over the devil.
negative The book does not live up to it’s promise. There is no revelation of truth about our heavenly father, or any insight into Satan’s lies. It is simply a polemic diatribe against Satan, with no concrete solutions to any of life’s problems.
Table 26: Generated examples for Polarity task using SynthesizRR.
Appendix I Data Preprocessing
Country of headquarters No. articles Domain
Antigua 2.6K antiguaobserver.com
Azerbaijan 70.7K trend.az
Bangladesh 28.2K thedailystar.net
Barbados 5.1K caribbean360.com
Brazil 930 thebraziltimes.com
China 10.7K chinadigitaltimes.net, china.org.cn
Colombia 22.9K colombiareports.com, insightcrime.org
Costa Rica 18.9K ticotimes.net
Cuba 1.6K escambray.cu
Cyprus 13.2K cyprus-mail.com, dailyforex.com
Czech Republic 1.2K praguepost.com
Egypt 43 thedailynewsegypt.com
Estonia 21.2K err.ee
Ghana 5.2K ghanabusinessnews.com, modernghana.com
Guyana 70.2K stabroeknews.com
Hong Kong 5.6K asiasentinel.com, actionforex.com, hku.hk
India 886.5K mid-day.com, financialexpress.com, livemint.com, hindustantimes.com, indianexpress.com, mangalorean.com, vccircle.com, deccanchronicle.com, afaqs.com, bollywoodhungama.com, medianewsline.com, orissadiary.com, morungexpress.com, countercurrents.org, businessworld.in, governancenow.com, koimoi.com, milligazette.com, dayafterindia.com, truthdive.com, newstodaynet.com, centralchronicle.com, dalje.com, rtn.asia, realbollywood.com, mutiny.in
Indonesia 2K thejakartaglobe.com
Iran 7.2K tehrantimes.com
Israel 60.4K jewishpress.com, ynetnews.com, palestinechronicle.com, 972mag.com, defense-update.com
Jamaica 96.6K jamaica-gleaner.com
Japan 2.1K japantoday.com
Kenya 158.8K capitalfm.co.ke, nation.co.ke, theeastafrican.co.ke, standardmedia.co.ke, kbc.co.ke, businessdailyafrica.com
Kuwait 16.2K arabtimesonline.com, kuwaittimes.net
Lebanon 4.9K yalibnan.com
Macau 3.4K macaudailytimes.com.mo
Malawi 2.8K maravipost.com
Malaysia 30.5K malaysiakini.com, freemalaysiatoday.com, theborneopost.com
Misc. Africa 51 african-bulletin.com
Misc. Asia 30.9K eurasiareview.com
Namibia 20.2K newera.com.na
Nepal 2.2K thehimalayantimes.com
Nigeria 336.5K thenationonlineng.net, vanguardngr.com, thisdaylive.com, codewit.com, sunnewsonline.com, businessdayonline.com, pmnewsnigeria.com
Pakistan 274.1K nation.com.pk, dawn.com, tribune.com.pk, pakobserver.net, app.com.pk, dailytimes.com.pk, thefrontierpost.com, pakistankakhudahafiz.com, thenews.com.pk, pak1stanfirst.com, pakwatan.com
Palestine 655 intifada-palestine.com, paltelegraph.com
Peru 4.6K livinginperu.com
Philippines 25.1K sunstar.com.ph, journal.com.ph, bworldonline.com, newsbytes.ph, mindanews.com, tribwekchron.com, philstar.com
Qatar 8.8K aljazeera.com, middle-east-online.com
Romania 13.3K zmescience.com
Saint Kitts and Nevis 4.6K thestkittsnevisobserver.com
Saudi Arabia 42.8K arabnews.com, saudigazette.com.sa
Singapore 112.4K straitstimes.com
Somalia 197 mareeg.com
Somaliland 4.7K somalilandpress.com
South Africa 22.9K itweb.co.za, memeburn.com, themediaonline.co.za, news24.com, iafrica.com, mybroadband.co.za
South Korea 22K koreatimes.co.kr, yonhapnews.co.kr
Sri Lanka 33.8K lankabusinessonline.com, onlanka.com, lankanewspapers.com, groundviews.org
Tanzania 7.6K thecitizen.co.tz
Thailand 11.2K pattayamail.com
Trinidad 3.2K trinidadexpress.com
Turkey 2.5K theminaretonline.com, nationalturk.com, melodika.net
Uganda 6.7K monitor.co.ug
United Arab Emirates 108.8K emirates247.com, gulfnews.com, ameinfo.com, meed.com, 7days.ae
Venezuela 3.9K venezuelanalysis.com
Zambia 7.4K lusakatimes.com
Zimbabwe 26.1K newsday.co.zw, nehandaradio.com, thezimbabwemail.com
Table 27: News domains from underrepresented countries in RealNews.
I.1 Datasets
Aside from ToI Headlines, we use the original datasets, randomly subsampling as mentioned in Table 1.
I.2 Corpora
  • RealNews: we use the article text field and download the data from https://github.com/rowanz/grover/tree/master/realnews.
  • RealNews/Regional is a subset of RealNews (Zellers et al., 2019). It includes 2.7M articles from non-US and non-EU websites. We manually check RealNews websites and identified 141 regional-news websites with headquarters in 56 non-US and non-EU countries: India, Pakistan, Nigeria, Philippines, etc. The complete list is mentioned in Table 27.
  • RealNews/India is further filtered to only include Indian news websites. We use only the “India” domains from Table 27.
  • RealNews/Dominant is the remaining 30.1M articles from 1063 news websites headquartered in 20 countries (of which over $75\%$ are US-based).
  • Products: We pull the data from https://nijianmo.github.io/amazon/index.html#complete-data and concatenate title and description.
  • CMU Movie Summary: Data is obtained from https://www.cs.cmu.edu/~ark/personas/, where we use the plot summaries file.
Appendix J Teacher and Student hyperparameters
J.1 Teacher LLM hyperparams
For LLaMa-2 Chat 13B, we use the implementation from HuggingFace: https://huggingface.co/TheBloke/Llama-2-13B-fp16 and run it at half-precision.
For Claude Instant-v1, we use Claude Instant v1.2: https://www.anthropic.com/news/releasing-claude-instant-1-2
We use a batch size of 1 for all generations as we have long contexts and encountered failures with higher batch sizes. We use nucleus sampling with top-p=0.9.
J.2 Student LM hyperparams
We use the same hyperparameters for DeBERTa-v3L and DistilBERT as (Yu et al., 2023a):
  • DistilBERT: Learning rate of 5e-5, gradient_accumulation_steps of 1, batch_size 32. We use the Adam optimizer with weight_decay of 1e-4 and epsilon of 1e-6. We use max_sequence_length of 512.
  • DeBERTa-v3L: Learning rate of 2e-5, gradient_accumulation_steps of 8, batch_size 4. We use the Adam optimizer with weight_decay of 1e-4 and epsilon of 1e-6. We use max_sequence_length of 512.
We train all students for 6 epochs. Following (Yu et al., 2023a), we use warmup for 6% of the training steps.
J.3 Oracle model hyperparams
To train the DeBERTa-v3-Large oracle model for Label Preservation, we use a grid search over 9 combinations: 3 learning rates {2e-5, 5e-5, 1e-4} by 3 batch-sizes {1, 4, 16} (with same graident accumulation). We train on 80% of the Gold training data and use the remaining 20% as validation.
J.4 Retriever
We use Contriever from HuggingFace library: https://huggingface.co/facebook/contriever.
We pass a batch-size of 512 for embedding.
Appendix K Computational budget
We run all our models on AWS Elastic Cloud Compute [^3] https://aws.amazon.com/ec2/ using 20 p3dn.24xlarge machines to call AWS cloud services, host the retrieval index and distill student models.
K.1 Information Retrieval
The corpora was embedded by us and the trivial was done using the Faiss library. [^4] https://faiss.ai/index.html We orchestrate 80 copies of Contriever using the Ray distributed framework [^5] https://docs.ray.io/en/latest/index.html to embed the RealNews and Products corpus in $\sim$3 hours each.
K.2 Dataset synthesis
In order to run LLaMa-2 Chat 13B and Claude Instant-v1, we invoke AWS Bedrock [^6] https://docs.aws.amazon.com/pdfs/bedrock/latest/APIReference/bedrock-api.pdf using the boto3 library [^7] https://boto3.amazonaws.com/v1/documentation/api/latest/index.html.
Generations were done at an AWS-account level RPM of 1600 and takes roughly 4 hours for a dataset of 8k rows.
K.3 Student distillation
Each DeBERTa-v3-Large student model trains for 1-3 hours on a single GPU on 8k rows. Each DistilBERT student model trains in 1 hour to generate the data-map for dataset catrography.
Appendix L Licensing
We use datasets that have been released in prior work with various open licenses. Specifically:
L.1 Datasets
L.2 Corpora